### v1

# Media Coverage of War Victims: Journalistic Biases in Reporting on Israel and Gaza

Abstract
October 7th 2023 marked the start of a war against Gaza, which is considered one of the most devastating wars in modern history and has led to a stark attitudinal divide within and between countries. To investigate the role of media bias in reporting on this asymmetrical warfare, we analyzed over 14,000 news articles published during the first year of war in three Western (The New York Times, BBC, CNN) and one non-Western English-language outlets (Al Jazeera English). Exploring the media narratives concerning Israeli and Palestinian victims experiencing hardship, we found three systematic biases in Western media. 1) Compared to Palestinian victims, represented mainly as undifferentiated collectives, Israeli victims were more likely to be portrayed as identifiable individual human beings. 2) Despite the striking difference in all forms of hardship (casualties, displacement, etc.), Western journalists created a false balance, equating Israeli and Palestinian suffering, by persistently referring back to the 7th of October massacre, even in the absence of new events involving Israeli victims. 3) When reporting on numbers of Palestinian (vs. Israeli) victims, journalists used language that casts doubt about the credibility of the information and the reputation of the source providing it, thereby selectively undermining the reader’s trust in the information regarding Palestinian suffering. Together, our analysis reveals a series of systematic journalistic biases in high-profile Western media that are absent or greatly reduced in Al Jazeera.
Introduction
More than other geopolitical events, the ongoing war on Gaza has produced an attitudinal divide between Western and Non-Western countries, which is reflected at the level of governmental decision making, support for UN resolutions, and public opinion. In the present research we investigate the role of media bias that may reflect, and contribute to, this attitudinal gap.
The ongoing U.S.-backed war by Israel against Gaza began on October 7th, following an attack by Hamas’ military wing on Israel. The attack resulted in over 1,200 deaths, including 38 children, the taking of approximately 250 hostages, and nearly 5,500 injured [1]. The attack is considered the deadliest attack in Israel’s history. This was followed by an extensive war of Israel against Gaza that - already by the end of the first year - had produced an unprecedented humanitarian crisis [2], the destruction of over half of Gaza’s homes [3], the devastation of the school, university, and health systems, extreme deprivation, and the forcible displacement of almost the entire population [1]. After one year of war, the death toll among Palestinians in Gaza amounted to over 42,000 people [4], the majority of whom women and children, although the true death toll may exceed 186,000 once indirect deaths are taken into account [5]. Over 97,000 people were injured [4], one fourth of whom with life-changing injuries such as amputations [6]. Different from other conflicts (such as Ukraine), the population is trapped in Gaza due to the 17-year-old air, sea and land blockade. The war has been particularly dire on children. In the first twelve months alone, over 14,000 had been killed [7], at least 17,000 were unaccompanied [8], and 90% suffered acute malnutrition [9]. According to Save the Children, more than 10 children each day had one or both legs amputated [10], and by the time of the writing, the humanitarian crisis has tragically deteriorated and multiple organizations, including the United Nations [11], Human Rights Watch [12], and Amnesty International [13], have declared that a genocide is taking place.
This war is also considered the deadliest conflict in history for aid workers with over 300 killed and for journalists with 128 killed within the first 12 months [14] and many others missing, injured or arrested [14, 15]. As a unique case in the history of journalism, news reports have relied almost exclusively on local journalists, given that international reporters were prevented from reaching the war zone, except for sporadic visits as embedded journalists under the strict control of the Israeli military [16]. Since the beginning of the war, foreign media such as The New York Times, BBC, CNN, ABC, NBC, and Fox News, have very limited access to Gaza, can operate only in the presence of Israeli soldiers, and “have to submit all materials and footage to the IDF for review before publication[17]. Thus, international reporters are prevented from baring witness to the tragedy unfolding in Gaza. Beyond the impact in the Middle East, the war also had global effects on intergroup relations, as reflected in a considerable increase in both anti-Jewish and anti-Muslim hate crimes since October 2023 [18, 19].
Against this backdrop, the present study examines how such a humanitarian disaster and information bottleneck are reflected in international media coverage. The current analysis, informed by psychological research and theory, focuses on subtle media biases in war journalism. Specifically, we examine three critical facets: (1) (De-)humanizing reporting, assessing to what extent victims are portrayed as individual human beings or as undifferentiated collectives. Thus, we look at the balance between group-level reports and individualized stories, which are more likely to evoke empathy. (2) Quantifying the human costs of war, exploring whether media reports present numbers in a meaningful and proportional manner, accurately reflecting the distribution of victims (and casualties more specifically) over the 12-month period examined. (3) Reporting on child victims, focusing on how children are individualized and featured in media coverage, as well as how their suffering is quantified.
Results
In this study, we analyze over 14,000 news articles published between October 7, 2023, and October 7, 2024, from The New York Times (NYT), BBC, CNN, and Al Jazeera English (AJE). We focus solely on the textual content, deliberately excluding other media formats like videos, podcasts and transcripts to maintain consistency in the analysis. For a detailed breakdown of the data composition and collection methodology, refer to the Materials and Methods’ Data subsection. As we were interested in subtle biases, we purposefully focused on highly professional news outlets, while excluding tabloids or other popular news sources where blatant partisanship could be expected.
We start our analysis by examining the data landscape presented in Figure 1, which illustrates the total article counts by source (Figure 1a) alongside the temporal distribution of article counts across sources (Figure 1b). The data reveal that NYT published the most articles, followed by AJE and CNN, with BBC contributing the least. The temporal distribution highlights significant spikes in response to key events, such as the October 7th attacks, the World Central Kitchen massacre in April 2024, attacks on Israel in both April 2024 and Sep 2024, as well as the Israeli attacks on schools and refugee buildings during Sep 2024; for a full list of the events and their exact dates please refer to Supplementary Tables 1 and 2. Furthermore, each outlet exhibited distinct sensitivity patterns: NYT consistently responded more than the others. CNN seems to adopt a similar overall publishing profile to that of NYT with publication spikes after relevant events, which, however, abate faster than NYT’s, suggesting they do not dwell on the event as long as NYT. The BBC maintained the lowest publishing profile and demonstrated a selective sensitivity to events, spiking its publications for some events, while remaining unresponsive to others. AJE, by contrast, maintained a relatively high publication rate, which remained comparably steady even after major events.
With this data, we conducted our key analyses. We investigated how media outlets handled (de-)humanizing reporting, quantified the human costs of war, and reported on child victims. This allowed us to uncover nuanced media biases across different sources.
(De-)humanizing the Victims: Individualized vs Category-based Reporting
War journalism has been criticized for its primary focus on frontline reporting, often neglecting the human consequences for the civilian population, and for children in particular [20]. The media may either quantitatively give little space to civilian victims or report on them in a way that conceals their full human qualities (dehumanization), including their intellect, their culture, their morality, their feelings, and their individuality. One way in which journalists may downplay the civilian experience of war is by reporting group-level experiences rather than individual stories. Social scientists such as Slovic [21] and Schelling [22] have long observed that people often remain indifferent to mass atrocities such as war, mass murder, and genocide, yet are profoundly moved by individual stories. This psychological tendency is captured by what scholars call identifiable victim effect [23]: people feel more compassion toward single victims and are more likely to take action (for instance, through donations) when text or images describe the suffering of a single individual rather than that of larger groups. Compassion quickly fades as the number of victims increases [21], with the exception of very cohesive and interconnected small groups such as families [24]. For instance, in a classical study by Kogut and Ritov [25], a single child in need of a costly life-saving treatment elicited greater distress and received greater donations than a group of sick children with identical needs. This phenomenon is driven by different mechanisms, most notably that identifiable victims attract greater attention and evoke greater emotional reactions (such as sympathy or distress) than statistical victims or large groups [21]. Also, by providing personalized information (such as age or name), as is typically done for single identifiable victims, narratives tend to be more vivid, which in turn induces a sense of familiarity and closeness. Unsurprisingly, the identifiable victim effect is more powerful when the victim is not responsible for their plight [23]. This is particularly relevant when reporting on children who, by definition, cannot be held responsible for an armed conflict or for the causes that led up to it. As such, by examining both the degree of individualization in war reporting and the specific attention devoted to children, our analysis seeks to uncover whether media narratives amplify or diminish the humanity of those most powerless in the face of war.
We began our analysis by identifying all instances in which individuals or groups were mentioned in news articles across our four media sources. To do this, we employed a large language model (LLM) to extract every reference to civilians, whether individual or collective, and classify each instance as pertaining to either the Palestinian or Israeli side. Specifically, we instructed the model to identify individualized civilian stories that highlight personal experiences of hardship, such as: “Noralin ‘Nataly’ Babadilla, 60, was visiting Kibbutz Nirim, to celebrate the community’s 70th anniversary. Her husband was murdered….” [26]. These types of narratives humanize the conflict by centering on a specific, emotionally resonant character. In addition, we asked the model to categorize the most severe type of hardship being described, enabling us to analyze not just whether individuals are mentioned, but also how their suffering is framed. Group-level references, which are typically broader and less personalized, included generalized hardships, such as “With many Israelis traumatized by the attack…” [27], statistical accounts, such as “More than 40,738 people have been killed in Gaza” [28], and event descriptions, like “Gazans are dying of starvation” [29]. A detailed explanation of how these classifications were extracted and validated is provided in Supplementary Note 1.
Next, we examined the extent to which each side is represented through individualized versus group narratives. As shown in Figure 2a, Western media outlets, BBC, CNN, and NYT, consistently featured higher ratios of individualized to group mentions for the Israeli compared to the Palestinian side. This asymmetry was most pronounced in NYT, where Israeli individuals were referenced once for every six group mentions (ratio = 0.16), whereas Palestinians were individualized only once for every twelve group mentions (ratio = 0.08). Similar patterns emerged for BBC (0.20 vs. 0.12) and CNN (0.13 vs. 0.09), though the disparity was less stark (see Supplementary Table 3). In contrast, AJE showed a more balanced pattern, with nearly equal ratios for both sides (0.09 for Palestinians vs. 0.07 for Israelis). This suggests that Israeli individuals were more frequently singled out in Western media coverage, while Palestinians were more often portrayed as part of an undifferentiated collective. We formalized this asymmetry using a Generalized Estimating Equations (GEE) regression analysis (see Materials and Methods, subsection “Testing Individualization Bias with GEE Models”), which confirmed that Western outlets were significantly more likely to individualize the Israeli side in their reporting. Specifically, we find that the odds of a reported story being individualized are 13.6% lower for Palestinian instances compared to Israeli ones in the case of BBC, 27.0% lower in the case of CNN, and 38.9% lower in the case of NYT. No statistically significant difference was observed for AJE.
Imbalance in Casualty Reporting
We then turned our attention to the content of the stories themselves, categorizing each instance based on the most severe type of hardship described. These included categories such as ‘casualties’, ‘displacement and refugees’, ‘imprisonment and detention’, among others. For a complete list of hardship categories and their distribution across media outlets, see Supplementary Figure 1 and Supplementary Note 1. Among these, we opted to concentrate on stories labeled as “casualties” (i.e., deaths and injuries), as this category was the most frequent across all sources, relevant to both sides, and supported by reliable statistics from international organizations such as the UN, allowing for systematic comparison.
When examining the ratio of Palestinian to Israeli casualty-related stories, as shown on the left side of Figure 2b, a stark contrast was revealed between AJE and Western outlets. AJE published roughly eight Palestinian casualty stories for every Israeli story, while BBC, CNN, and NYT maintained a close to one-to-one ratio. NYT even showed a slight tilt in the opposite direction, reporting eight Palestinian stories for every ten Israeli ones. At first glance, this might suggest a pro-Palestinian bias in AJE’s reporting. However, when we compared these ratios to the actual casualty figures, used here as a crude baseline for contextualizing media attention, a more complex picture emerged. As shown on the right side of Figure 2b, for every Israeli casualty, there were roughly 20 Palestinian casualties by the end of the first year of war. This stark disparity in actual casualties reflects the marked power differential between the two groups, and contrasts with the seemingly “balanced” reporting in Western media. In theory, if all events were covered equally and without selection bias, one would expect a proportional discrepancy in the number of stories reflecting the underlying reality. Instead, the reporting style of Western media and, to a lesser degree of AJE, resembles what is known as “false balance”, namely the tendency to create the impression that two sides are equivalent when, in reality, they are not (as will be explained in more detail in the discussion section).
Sustaining Parity in the Face of Discrepancy
This led us to a critical line of inquiry: How do Western media outlets sustain parity in reporting casualty-related stories between both sides when the actual number of casualties is so drastically imbalanced? To investigate this, we analyzed the content of Israeli casualty-related stories published by Western media (generated using Prompt 2 in Supplementary Note 1). Strikingly, we found that 92% of these stories were related to the events of October 7, as shown in Figure 2c. This concentration persisted even in stories published several months after the attack. In other words, a single moment in time, the initial event, was repeatedly covered to sustain a consistent volume of Israeli casualty narratives, thereby flattening the temporal and numerical asymmetry in actual deaths.
To explore this pattern further, we examined how Western media reported on the Israeli side during moments of significant Palestinian civilian casualties. We began by testing whether spikes in Palestinian deaths were associated with renewed publication of October 7–related Israeli stories. Indeed, a strong positive correlation emerged (Pearson’s r = 0.8; see Supplementary Figure 2), suggesting that Israeli narratives were often recycled precisely when Palestinian deaths were peaking. To better understand how this dynamic played out in practice, we selected four equidistant dates of high Palestinian casualties across the year and reviewed the stories that appeared on those dates and the day after.
Across these cases, a clear pattern emerged: while Palestinian tragedies were reported, Israeli stories received more narrative depth, emotional detail, and prominence, even on days when no new Israeli casualties occurred. For instance, on 1 December 2023, nearly 184 Palestinians were killed [30] as Israel resumed its assault following a ceasefire, yet NYT ran stories quoting hostage families: “Mentally, she’s definitely doing better. She’s more engaged with people and becoming more independent.”[31], while CNN reminded readers that “around 240 people, from infants to octogenarians, were taken hostage during Hamas’ attack on October 7” [32]. Across two days, BBC, CNN, and NYT published 19, 21, and 24 Israeli stories respectively, versus just 17 Palestinian stories combined. A similar imbalance appeared during the “Flour Massacre” on 29 February 2024, when 117 Palestinians were killed and 750 injured [33] while receiving food aid. Despite the scale of the tragedy, Western outlets again devoted substantial space to Israeli narratives, collectively running 10 Israeli-focused stories versus 16 on Palestinians, a ratio that muted the scale of the Palestinian tragedy. CNN even described the incident as the “worst tragedy” yet, but in the same news cycle published an opinion piece urging readers: “When you kiss them goodnight — please think of Kfir and Ariel Bibas. Think of their mother and father” [34].
The third date explored was 8 June 2024, which recorded the highest number of Palestinian casualties among the dates we examined: 274 killed and nearly 700 injured [35]. Given this extraordinary scale, and the tentative shift toward balance observed on 29 February, one might have expected coverage to tilt more heavily toward Palestinians. This was not the case, as the date also coincided with a major Israeli development: the rescue of four hostages from Gaza’s Nusseirat area. Coverage of the rescues was celebratory,“Saturday was an emotional and happy day for the state of Israel and the IDF” [36], while Palestinian casualties were reported with skepticism: “CNN has no way of verifying casualty numbers reported by Palestinian officials in Gaza” [36]. Across that day and the next, BBC, CNN, and NYT ran a total of 40 Israeli stories, more than double the 19 on Palestinians, raising the possibility that when Israeli and Palestinian events coincided, Israeli narratives overshadowed even the deadliest Palestinian tragedies.
By the fourth date explored, 10 September 2024, coverage appeared to move back toward a more balanced pattern. An Israeli strike on a humanitarian zone killed 19 Palestinians [37], and both BBC and CNN reported only on Palestinian casualties. The New York Times still ran hostage-focused stories, re-emphasizing their suffering: “the hostages suffered from significant malnutrition, severe weight loss and long-term physical neglect” [38], yet overall published more on Palestinians than Israelis (7 vs 4). This suggested that, with time, the narrative might have been shifting toward greater balance. However, the 8 June example, where Israeli and Palestinian events coincided, raised the question of whether this balance only held when Israeli developments were absent. To probe this further, we examined two additional dates where both sides experienced major events: 15 January 2024, marking 100 days since October 7, and 1 September 2024, following the recovery of six Israeli hostage bodies [39].
On 15 January 2024, Israeli airstrikes on Gaza’s Al Thalatheni Street killed 22 Palestinians and injured many more [40]. Yet the day was largely framed in Western media as the “100 days since October 7” milestone. Coverage focused heavily on hostage testimonies and recollections, such as NYT’s headline “They thought they knew death but that didn’t prepare them for Oct 7” [41] and personal accounts like “Israeli teenager recounts her time as a hostage in Gaza” [42]. BBC emphasized military perspectives with stories such as “They were Israel’s eyes on the border but their Hamas warnings went unheard” [43]. Palestinian casualties, by contrast, received comparatively little attention. Across two days, BBC, CNN, and New York Times ran 13, 10, and 12 Israeli stories, compared to just 6 on Palestinians combined, with NYT publishing none at all. This imbalance suggests that symbolic Israeli milestones may have drawn coverage away from concurrent Palestinian losses.
On 1 September 2024, an Israeli strike on a school-turned-refugee camp killed 11 Palestinians, including several asleep in their tents [44]. Yet coverage that day was dominated by the recovery of six Israeli hostage bodies [39] the previous evening. Western outlets devoted extensive space to personal tributes: the BBC described one victim as “among the innocents brutally attacked while attending a music festival for peace… He had just turned 23. He planned to travel the world” [28]. NYT highlighted family memories: “Hersh Goldberg-Polin loved soccer and music. He was curious, respectful and passionate about geography and travel” [45]. CNN portrayed him as “one of the most recognizable faces of the enduring hostage crisis… Banners and murals demanding his return were often displayed in Jerusalem and around the world” [46]. In contrast, Palestinian victims were mentioned only briefly. Across 1–2 September, BBC, CNN, and New York Times published 86, 38, and 57 Israeli stories, compared to just 5 Palestinian ones combined, with CNN publishing none. Here again, the overlap of major events may help explain why Israeli narratives overshadowed Palestinian casualties, despite the ongoing humanitarian toll.
Taken together, these cases show that although Western media did cover major events affecting both Palestinians and Israelis, at times even including vivid depictions of Gaza’s devastation, such as describing Al-Ahli Hospital as “like a horror movie” [47] or reporting that “children [were] torn apart and scattered in the streets” [36], the overall volume, depth, and emotional richness of coverage skewed heavily toward the Israeli side. Coverage appeared more balanced only when no major Israeli events occurred. When significant events on both sides coincided, however, Israeli narratives were amplified to such a degree that they overshadowed even the most severe Palestinian tragedies (see Supplementary Table 4). Israeli stories were also far more frequently personalized, with names, ages, aspirations, and family histories highlighted [45, 46, 28], and often revisited the same individuals over time. This suggests that the imbalance was not merely the result of coverage decay but was actively reinforced whenever Israeli stories could be interwoven with Palestinian casualty events, thereby sustaining a structural asymmetry in how loss and victimhood were represented.
A more detailed explanation of our date-selection procedure and findings, including additional quotes, is provided in Supplementary Note 2. The number of Israeli stories, related to October 7, published by each outlet on the dates examined can be found in Supplementary Table 4.
Quantifying Empathy-Related Narrative Elements
To assess whether the qualitative imbalance we observed persisted across the full year of coverage, we quantified differences in narrative elements that can influence reader empathy, using the validated Human Empathy and Narrative Taxonomy (HEART) framework [48]. Specifically, we applied two HEART metrics: (i) Vividness of Emotions and (ii) Plot Volume (the main predictors of perceived empathy in the HEART model) to all hardship-related individualized stories in our corpus. While these metrics are not direct measures of empathy, prior research shows they are key stylistic features that contribute to how much empathy a narrative may elicit. Scores for each metric were first calculated at the story level and then aggregated by side within each article, given that articles often addressed both sides. For each metric, we then calculated a difference score by subtracting the Israeli aggregate score from the Palestinian one, with positive values indicating that these empathy-related narrative elements were more prominent in Palestinian stories, and negative values indicating the reverse. For full details on the scoring procedure, see “Measuring Narrative Elements Associated with Empathy in News Texts” in the Materials and Methods.
Weekly averages of the difference scores for each outlet are shown in Figure 2d (Vividness of Emotions) and Supplementary Figure 3 (Plot Volume). In these plots, values above 0 indicate that, on average, stories published during that period were more vivid or detailed for the Palestinian side; values below 0 indicate the reverse, and 0 represents parity.
AJE consistently leaned toward more vivid Palestinian narratives, though without dramatic fluctuations, mostly hovering between 0 and 2. In contrast, all three Western outlets showed two clear spikes toward the Israeli side: one in the final two weeks of November 2023, coinciding with the hostages–prisoners swap, and another in the last two weeks of August 2024, when hostage negotiations collapsed and six Israeli hostages were found dead in Gaza. In both cases, the spike in vividness for Israeli stories exceeded even AJE’s highest levels for Palestinians.
The Plot Volume results show the same pattern, including the two spikes, with Israeli-focused stories in those periods receiving even higher relative scores than in the Vividness of Emotions metric. At the same time, it is worth noting that, on average, stories about Palestinians were written with greater emotional vividness across outlets (albeit less so for NYT), which is not surprising given the scale of Palestinian suffering. What is striking, however, is that even outlets such as BBC and CNN, while often writing about Palestinians in a more emotional way, still devoted disproportionate attention to Israeli narratives during key hostage-related events. Taken together, these findings suggest that the asymmetry lies less in the tone itself than in the balance of coverage: Western outlets often create a false equivalence by giving equal or greater narrative weight to Israeli experiences, despite the vastly different realities of suffering on the ground.
False Balance in the Coverage of Child Victims
Given the high vulnerability of children and the fact that they cannot be held responsible for armed conflict nor for the election of leaders who promote armed conflicts, it is not surprising that children have a special status in international humanitarian law (among which the UN Convention on the Rights of the Child [49]). Similarly, lay people value the life of the young higher than that of adults [50], feel a greater moral obligation to help children [51], and are more likely to assist them [52, 53]. Children (including older children) are prioritized in helping decisions because they evoke greater nurturing concern and compassion [54], because they are generally not held responsible for their fate, and because they have, on average, a greater number of years ahead, which makes helping them a rational choice. As such, it is essential to understand how much attention the different media sources pay to child victims of war. We therefore extended our analyses of individualized reporting to specifically track mentions of children.
On average, the four media outlets dedicated roughly 46.5% of their individualized stories to children (AJE 47%, BBC 45%, CNN 50%, NYT 44%; Supplementary Table 5a). Taken at face value, this looks like responsible reporting: the media gave children more prominence than their actual share of casualties (about 32%), which seems appropriate given children’s heightened vulnerability and moral salience. However, when the data are broken down by side, the picture changes. Palestinian children do take up a commendable share of stories—roughly in line with their 34% share of deaths—but Israeli children receive far more attention than their casualty numbers would suggest. Although they make up only 3% of Israeli deaths, their stories account for as much as 44% of child-related individualized reporting in CNN, with similarly elevated shares in BBC (36%) and NYT (40%). This disproportionate emphasis mirrors the false balance documented in our earlier analyses, where the relative weight given to Israeli and Palestinian suffering diverged sharply from the underlying reality.
The imbalance becomes even clearer when looking at the distribution of child-related stories within each outlet (Supplementary Table 5b). If each child’s life were given equal weight, nearly all individualized stories about children (over 99%) would focus on Palestinian victims. Instead, Western outlets devoted a strikingly large share of their child coverage to Israeli children: BBC 44%, CNN 43%, and NYT 51%. This pattern reflects a form of false balance: by the end of the first year of war, 38 Israeli [55] and 14,000 Palestinian children [7] had been killed, a disparity likely even greater for other hardships such as injuries, displacement, and famine. If casualty numbers alone were used as a baseline, individualized stories on Israeli children would be expected to account for less than 1% of all child references. AJE came closest to reflecting the underlying disparity, with 91% of its child stories about Palestinians. The most extreme deviation is observed in the New York Times, where Israeli and Palestinian children were given nearly equal attention, despite the enormous difference in the actual number of victims.
Quantifying the Human Cost of War: the Importance of Numbers
The reporting of victim numbers in news articles is crucial for conveying the magnitude of casualties and the impact of violence on communities. Quantitative data not only provides readers with a clear understanding of the scale of incidents but also shapes public perception and policy responses to crime and victimization. Research indicates that the inclusion of victim statistics can significantly influence how audiences interpret the severity of an issue, thereby fostering a sense of urgency and empathy towards affected populations [56]. For instance, when media outlets report on violent crimes, the number of victims can serve as a stark reminder of the human cost associated with such events, prompting discussions on prevention and intervention strategies [57]. Furthermore, the reporting of victim numbers can also reflect broader societal attitudes towards certain groups, as disparities in reporting are likely to lead to the devaluation of victims from marginalized communities [58]. This selective emphasis on victim statistics not only affects public discourse but can also influence legislative and law enforcement priorities, underscoring the media’s role in shaping narratives around crime and victimhood [59]. Thus, accurate and comprehensive reporting of victim numbers is essential for fostering informed public dialogue and effective policy-making.
Recognizing the importance of reporting numbers in conflict scenarios, we employed an LLM to analyze all articles and extract instances containing numerical references to civilians being harmed in any capacity, such as “100 civilians were killed in that strike.” For clarity and convenience, these instances will be referred to as Civilian Victim Numbers (CVNs). A detailed explanation of how this information was extracted and validated can be found in Supplementary Note 3. We begin by examining the total number of articles containing at least one CVN. Figure 3a presents a bar plot displaying the number of articles for each of the four media sources. Each source is represented by three bars: one showing the total number of articles (grey), one for articles mentioning CVNs related to the Palestinian side (green) and one for the Israeli side (blue). Notable differences in reporting styles emerge when comparing AJE with the Western news outlets. AJE reports twice as many articles with Palestinian CVNs than Israeli, while BBC and CNN report on both sides almost equally. Furthermore, NYT presents an interesting case: although the absolute number of its articles mentioning Palestinian CVNs (2,705) is comparable to AJE’s (2,847), the proportion is markedly lower (46% for NYT versus 74% for AJE). At the same time, NYT reports more on Israeli CVNs in absolute terms than any other outlet, whereas the BBC devotes the highest proportion of its coverage to Israeli CVNs (59%).
Given the significant disparity in civilian casualties reported by October 7, 2024 (over 41,689 Palestinians versus 1,200 Israelis killed, and 96,625 Palestinians versus 5,432 Israelis injured, as documented by the United Nations [4]), this imbalance should logically be reflected in media reporting, with greater attention directed toward the side experiencing more severe and prolonged impacts. To evaluate this, we assessed the proportion of CVNs specifically related to casualties (for a breakdown of the different types of CVNs, refer to Supplementary Table 6). Our findings confirm that the majority of CVNs were indeed linked to casualty-related statistics. Moreover, we observe that, across all outlets, Palestinian CVNs outnumber Israeli CVNs. However, in Western media specifically, Israeli CVNs still amount to more than half the number of Palestinian CVNs, despite the vast disparity in actual casualties. Once again, this points to a form of false balance, where Israeli suffering is elevated to a level of prominence in coverage that does not reflect the underlying reality.
False Balance in Coverage: A Baseline Comparison of Casualty-Related Numbers
To assess whether the observed imbalance is limited to the early stages of the war following the October 7 attack, we analyzed the aggregate article counts from each news source over time. Supplementary Figure 4 illustrates the difference in the percentage of articles mentioning Palestinian casualty-related CVNs versus Israeli casualty-related CVNs. Data points above zero indicate more articles reporting Palestinian CVNs than Israeli ones, and data points at zero indicate equal reporting for both sides. The data reveal that AJE consistently reported more articles on Palestinian casualties, with an average percentage difference of approximately 41%, whereas BBC, CNN, and NYT hovered closer to zero, averaging 10%, 11%, and 14%, respectively. Given the substantial disparity in casualty numbers between the two sides, a higher proportion of Palestinian CVNs is expected. Moreover, while Israeli casualties were concentrated largely around October 7 and its immediate aftermath, Palestinian casualties were sustained throughout the year, which would further suggest more frequent reporting on Palestinians to reflect this temporal distribution. However, the question arises: how large should that difference be? Does a line far above zero indicate bias toward Palestinians, while one close to zero would reflect bias toward Israelis? Does AJE’s reporting reflect fairness, while other outlets lean toward bias, or is the opposite true? Determining at what point the percentage difference transitions from fair reporting to bias requires establishing a baseline to objectively evaluate the reporting styles of the four media sources.
To this end, we propose a baseline model to predict the expected number of mentions of casualty-related CVNs in news outlets based on weekly casualty numbers reported by the United Nations Office for Coordination of Humanitarian Affairs (OCHA) during the first year of the conflict [4]. The model calculates the expected share of mentions for Palestine and Israel, relative to their casualty numbers. This aligns with findings by Miller and Albert (2015), who observed that media coverage is strongly influenced by casualty figures, following the adage “If it bleeds, it leads,” wherein fatal incidents dominate narratives and amplify their emotional and psychological impact [60]. The model also accounts for sudden increases in media coverage, or “spikes,” caused by sharp rises in casualties. These spikes, reflecting significant shifts in violence, temporarily alter media dynamics, as shown by previous research on media reactions to crises [61, 62]. Additionally, the model incorporates a decay factor to capture the natural decline in coverage over time. This decay, often referred to as “agenda-decay,” is consistent with findings by Pfefferbaum et al. (2014) and Downing et al. (2004), which highlight that media attention diminishes without continuous stimuli, reflecting the temporal dynamics of waning public interest as the news cycle progresses [63, 64]. These components enable the model to account for the fluctuating patterns of media attention throughout the conflict. For a detailed description, kindly refer to the subsection “Baseline Model of Expected Weekly Casualty Reporting” in Materials and Methods.
The comparison between the expected number of casualty-related CVN mentions, based on the baseline model, and the actual number of casualty-related CVN mentions for each side is visualized over time for each news source in Supplementary Figure 5. A consistent trend emerges across all outlets: Israeli CVNs appear to be overreported, while Palestinian CVNs seem to be underreported. However, comparing the magnitude of these gaps, between expected and actual mentions, across different news sources is challenging. Just how much overreporting or underreporting, and thus potential bias, exists relative to the baseline? To address this, we calculate the weekly proportion of expected mentions allocated to each side based on the baseline model for each news outlet.
For example, suppose that in a given week $t$, the baseline model predicts that, in news source $n$, 90% of all casualty-related CVN mentions should concern Palestinians and 10% should concern Israelis, based on the actual number of deaths that week. In reality, however, news source $n$ might report 60% of its casualty-related CVN mentions about Palestinians and 40% about Israelis. This would indicate a 30-percentage-point shortfall in Palestinian coverage ($\Delta P_{n}^{t}=+30\%$) and an equivalent overrepresentation of Israeli coverage for that week ($\Delta I_{n}^{t}=-30\%$). This ensures the differences are symmetric in absolute terms, reflecting the same magnitude of deviation for both sides. We display the Palestinian gap ($\Delta P_{n}^{t}$) to visualize the discrepancy between actual reporting and the baseline expectation, allowing us to assess the bias of each news outlet. These results are illustrated in Figure 3b. The figure reveals that NYT, CNN, and BBC exhibit a greater bias toward Israel, with the US-based sources (NYT and CNN) frequently interchanging over time. BBC stands out as the most biased news source overall, with an average gap of approximately 31%. In contrast, AJE consistently demonstrates the least bias, maintaining an average gap of just 13% over time in favor of Israel.
Selective Citation of Sources
Our analysis thus far has focused on whether one side’s CVNs are being under-reported, operating under the assumption that these numbers are fact-checked and accurate, thereby granting media sources the benefit of the doubt. However, the reporting practices of mainstream media often reveal biases that can profoundly influence public perception and understanding of events. One prominent aspect of this bias is the selective citation of sources and facts, which serves to reinforce particular narratives while marginalizing alternative viewpoints. Research indicates that such selective reporting can create a skewed representation of reality, as news sources may prioritize information that aligns with their ideological leanings, thereby engaging in what has been termed “partisan coverage filtering” [65]. Furthermore, the phenomenon of confirmation bias suggests that media organizations may favor facts that support their existing narratives, while casting doubt on those that do not [66, 67]. This can manifest in the use of language that introduces uncertainty around facts that contradict the media’s preferred storyline, thereby shaping audience perceptions in a way that aligns with the outlet’s biases [68, 69]. The implications of these practices are profound, as they not only distort the truth but also contribute to a polarized media landscape where audiences are increasingly exposed to biased information that aligns with their pre-existing beliefs [70, 71].
Building on the previous observations, we conducted an analysis to investigate the possibility of selective citation of sources for CVNs. We utilized an LLM to evaluate each sentence containing a CVN, prompting it to determine whether a source was cited (for details, see Supplementary Note 4). Subsequently, we examined the frequency of cited sources for CVNs on each side and compared patterns across different news outlets (examples of the most popular ways to cite sources for each side can be found in Supplementary Tables 7 and 8). The results, detailed in Supplementary Table 9, highlight distinct differences in reporting practices. AJE demonstrated a balanced approach, citing sources with almost equal frequency for both Palestinian and Israeli CVNs, indicating a consistent standard for source attribution. In contrast, Western news outlets were significantly more prone to cite sources when reporting on Palestinian CVNs than Israeli ones. U.S.-based outlets, in particular, referenced sources for Palestinian data nearly twice as often as for Israeli figures. BBC, on the other hand, reported Israeli CVNs with the least source attribution, at only 16.9% of instances, while attributing sources to Palestinian CVNs three times more frequently. These disparities raise questions about potential biases in validating casualty figures based on the context.
Creating Skepticism
Next, we analyzed how referenced CVNs were phrased to determine whether they were presented in a “casting doubt” style (details can be found in Supplementary note 5). Using an LLM, we examined full sentences containing CVNs for phrases implying skepticism or questioning the source or the reported numbers. These doubt-casting phrases were categorized into two main types: (1) Source Doubting, i.e. phrases that question the credibility of the reporting source, thereby indirectly doubting the reported numbers, such as: “Israel’s relentless airstrikes have killed more than 8,000 people, according to the Hamas-run health ministry[72] and “The health ministry in Gaza, which is controlled by Hamas, estimates that 11,000 civilians have been killed there over the last month,” [73] and (2) Uncertainty in Numbers, i.e. phrases that cast doubt specifically on the reported numbers, such as: “…dozens of civilians were reportedly killed when jets struck buildings…” [74]. For a complete list of doubt-casting phrases and their categories, see Supplementary Table 10. Notably, our analysis found no source-doubting phrases on Israeli-reported numbers, even after using a second independent LLM prompt designed specifically to detect such instances. This suggests that doubt-casting techniques are disproportionately applied to Palestinian-reported numbers. While this study does not aim to verify the accuracy of these numbers, its purpose is to evaluate, through numerical analysis, whether both sides are presented equally in news coverage. We explore this further below.
After identifying doubt-casting phrases, we analyzed their occurrences for each side (Palestine vs. Israel) across four news sources. The results, summarized in Figure 3c, reveal that BBC overwhelmingly leads in using doubt-casting phrases, with 915 CVN mentions, 98% of which target Palestinians—averaging over two CVNs doubted daily during the first year. This far exceeds CNN (145), NYT (131), and AJE (72). Common Source Doubting phrases like “Hamas-run health ministry” were heavily used by BBC, CNN, and NYT, whereas analogous expressions for Israel (such as “Knesset-run health ministry”) were absent. CNN and NYT also favored phrases like “Hamas-controlled enclave.” For Uncertainty in Numbers, terms such as “reportedly” and “claim” were predominantly used to describe Palestinian casualties by Western outlets, whereas AJE applied them more evenly across both sides (see Supplementary Tables 11–14 for detailed lists of phrases by outlet). Western sources attributed 71–90% of their cast-doubting phrases to doubting the sources of Palestinian CVNs, compared to 21% for AJE. Additionally, AJE was the only outlet to use Uncertainty in Numbers phrases in a balanced manner: 53% referred to Palestinians and 26% to Israelis, a proportion consistent with the fact that AJE published roughly twice as many articles about Palestinian CVNs as about Israeli ones. In contrast, the other outlets applied such language almost exclusively to Palestinians, with 3% or less referring to Israelis. These findings underscore in particular BBC’s significant reliance on doubt-casting phrases, particularly when reporting on Palestinian casualties, compared to other sources.
Children in the Narrative: Imbalances in CVN Reporting
To conclude this section of our analysis, we examined how frequently media outlets referenced children in their reporting of civilian victim numbers (CVNs) for both Palestinians and Israelis (Supplementary Figures 6 and 7, and Supplementary Table 15). To contextualize these proportions, we benchmark them in Supplementary Table 15 against the approximate share of children among civilian deaths on each side: around 5% for Israeli civilians on October 7 (and roughly 3% when including soldiers, based on public compilation) [75, 55], compared to 40–44% for Palestinian civilians over the first year of the war (October 7, 2023–October 7, 2024) [76, 77].
Relative to these baselines, Al Jazeera English (AJE) provides the most proportionate coverage of Palestinian children, with casualty-related child mentions at 29.3% (closest to the 40–44% reference) but reports the least on Israeli children, at only 2% compared to the 5% baseline. The New York Times (NYT) exhibits the reverse pattern: it comes closest to the Israeli baseline (4.2% vs. 5%) but falls furthest below the Palestinian reference (15.7% vs. 40–44%). BBC and CNN occupy intermediate positions, with 20.8% of casualty-related Palestinian child mentions for both, and 2.6–3.4% for Israeli children. Overall, these patterns highlight how outlets differ not only in their quantitative attention to children but also in how proportionally, or selectively, they mirror the actual distribution of child victims across the two populations.
Discussion
Three main conclusions can be drawn from our media analyses of over 14.000 articles published by four international news outlets during the first year of the 2023 war on Gaza.
First, unlike AJE, Western media (BBC, CNN, and NYT) tend to emphasize the individuality of Israeli victims experiencing hardship, while treating Palestinian victims more frequently as a collective. This is evident in the higher ratio of individualized to group-level reporting for the Israeli compared to the Palestinian side. The only outlet showing a more balanced pattern is AJE; here the prevalence of individualized (vs. group-level) narratives is very similar for the two sides. These findings suggest that even highly professional (and relatively “liberal”) Western media tend to downplay the individuality of Palestinian victims experiencing the traumatic (and often fatal) consequences of war, while acknowledging, to a greater extent, the personal suffering of Israeli victims. Based on previous literature on the identifiable victim effect [21], the Western reporting pattern is likely to elicit greater empathy and compassion for Israeli than for Palestinian victims. Although testing the impact of this reporting bias on the readership was beyond the aims of the present study, this appears to be a worthwhile endeavor for future research.
Looking at the empathic style of reporting, we see a more complex pattern. The two main indices of the Human Empathy and Narrative Taxonomy [48] show a small but consistent tendency of AJE journalists to report in greater detail (plot volume) and with more vivid emotions on individual Palestinian than Israeli victims. In contrast, Western media appear more reactive to specific events, showing modest pro-Palestinian spikes at some moments and stark pro-Israeli spikes in others. Taken together, these findings suggest that it is not as much the style of writing than the under-reporting of individual stories that contributes to the dehumanization of Palestinians in Western media.
The second systematic bias in Western media (and to a lesser degree in AJE) was the tendency to equate Israeli e Palestinian suffering despite the stark difference in all forms of hardship (casualties, displacement, arbitrary detention and kidnapping, destruction of the health, school, and university systems, etc.). In the absence of bias, news reports should reflect the actual gaps in human suffering. To give only one example, if Israeli child deaths are less than 1% and Palestinian children more than 99% of the total number of children killed during the first year of war, news reports would roughly reflect this discrepancy. In contrast to this assumption, we find a systematic deviation from actual discrepancies that emerges consistently across observations: (a) In Western media, 47% of individual narratives on children referred to Israeli children who constitutes less than 1% of the victims; (b) Including all age groups, the ratio of Palestinian to Israeli casualty-related individual stories was approximately 1:1 in Western media, despite an actual casualty ratio of 20:1. (c) Especially in Western media, Israeli Civilian Victim Numbers are over-reported and Palestinian Numbers under-reported compared to what would be expected based on the actual number of victims and the distribution over time. Together, these reporting patterns insinuate a false premise of symmetry, where, in reality, the entity and prevalence of human suffering is blatantly asymmetrical.
How did BBC, CCN, and NYT journalists manage to create the illusion of balance in the face of extreme disparities in military power and human suffering? The main strategy was to persistently refer back to the 7th of October massacre, even at a distance of months and in the absence of new events involving Israeli victims. Our analysis suggests that reiterations of October 7 were particularly pronounced during periods when significant events occurred on both sides, often overshadowing coverage of Palestinian suffering and diverting attention from the toll in Gaza. This strategy, which contradicts the common “agenda decay” phenomenon, not only creates an illusion of balance, but also reiterates the idea that the 7th of October was the starting point and “cause” of the subsequent events.
This reporting pattern resembles what is generally referred to as “false balance” or “false equivalence”. False balance occurs when equal space or attention (e.g. airtime) is given to two sides that, in reality, differ in validity, such as climate skeptics and climate scientists [78] or pro- and anti-vaccination proponents [79, 69]. Extending this concept to war journalism, false balance can manifest when unequal parties in a conflict are portrayed as equally culpable or their actions as equally justified. Regarding victims, a false balanced perspective may portray different sides involved in a conflict as equally affected by the conflict. This can obscure asymmetries in power, suffering, and in violations of international law, leading audiences to perceive the conflict as an intractable clash of opposing but equivalent sides. As one of the main consequences of false balance is the production of uncertainty [80], it is plausible that, in the context of war journalism, falsely balanced portrayals reduce public recognition of victimization or aggression, thereby dampening calls for accountability and action. Whether this applies to the present case remains an open question for future research.
A third finding concerns the presumed reliability of the reported information and the trustworthiness of the source providing it. Unlike AJE, Western journalists were found to engage in strategies aimed at planting seeds of doubt in their readership when reporting on Palestinian victims. When quantifying the human cost of war for civilian victims, Western journalists tended to cite sources more frequently for Palestinian than for Israeli victims, as if Israeli (but not Palestinian) statistics can be taken for granted. This interpretation is further corroborated by a systematic and asymmetrical “casting doubt” style that questions the veridicality of data predominantly when referring to Palestinian victims (such as “reportedly”, “claims”, “according to the Hamas-run health ministry”, and the like). By discrediting sources in a subtle way, journalists undermine the reader’s trust in the information provided and in the reputation of the source.
Together, the present media analysis has identified multiple biases in Western media that, together, may have contributed to the global divide in public opinion and to the delay in public outrage in the Western world in front of the unfolding genocide. Somewhat surprisingly, we did not discover a complementary pro-Palestinian bias in AJE, which - with a few exceptions, like the underreporting of Israeli children CVNs - displayed a relatively balanced reporting style.
Limits and future directions
Despite its extensive coverage, the present study has a number of limitations that should be acknowledged and that could serve as a starting point for future research. First of all, we took a conservative approach and included only highly respected and relatively “liberal” media in our analysis. This has almost certainly led to an underestimate of actual media bias.
Second, our media analysis lacks a comparison to other conflict situations in order to understand whether the biases identified here are unique to Palestine and Israel. The Ukraine war may offer such a comparison standard, given the temporal proximity. Thus, future studies should not only broaden the scope to include more conservative media (such as Fox News or Daily Mail), but also extend the analysis to other conflict situations (such as Ukraine).
Another important future development regards the extension of the analysis beyond the first year of war. It is conceivable that the pro-Israeli / anti-Palestinian bias in Western media may have declined during the second year of the ongoing genocide and that this may have contributed to the greater awareness and the shift in public opinion in the West. Investigating whether the media biases identified here did, indeed, shape public opinion, and to what degree, remains another important issue to be addressed in future research. Only longitudinal or experimental methods will be able to address this question and allow inferences about the actual impact of the biases observed here on public opinion and collective action.
Materials and Methods
Data Collection
The New York Times (NYT), CNN, and Al Jazeera English (AJE) news articles were accessed via the LexisNexis Academic Web Services API (WSApi) [^1] LexisNexis NYU guide, https://guides.nyu.edu/lexisnexis-rest-api [^2] LexisNexis WSApi https://solutions.nexis.com/wsapi/, accessed through October 2024. These articles cover the period from October 7, 2023, to October 7, 2024. To ensure the inclusion of only highly relevant news articles, we excluded specialized content such as podcasts and audio transcripts. We also removed summary articles that condensed multiple topics into brief overviews. Additionally, we filtered the articles based on country and subject tag, retaining only those with content relevant to Palestine and Israel, specifically tagged with “Hamas,” “Israel,” “Gaza,” “West Bank,” and “Palestin.” (the term Palestin was used to account for the variations of the word, such as Palestinian(s) and Palestine).
For BBC news articles, which were not part of the LexisNexis dataset, we retrieved them directly from the BBC website. Similar to the LexisNexis filtering, we focused on articles categorized under tags pertinent to our research, including “Israel,” “Israel and the Palestinians,” “Gaza,” “Palestinian territories,” “Hamas,” “West Bank,” and “Israel-Gaza war.” Non-news media, such as live feeds and audio transcripts, were also excluded. Further, we used Google News Search to find relevant articles, restricting our query to return results that belong to the BBC website and we only kept articles with titles containing the following terms: “Hamas,” “Israel,” “Gaza,” “West Bank,” “IDF,” and “Palestine.” This process resulted in a dataset comprising 14,280 news articles, distributed as follows: 5,829 (40.8%) from NYT, 3,850 (27%) from AJE, 2,543 (17.8%) from CNN, and 2,058 (14.4%) from the BBC.
Testing Individualization Bias with GEE Models
To test whether one side was more likely to be individualized in news coverage, we conducted a series of Generalized Estimating Equations (GEE) logistic regression analyses. This approach accounts for potential correlation among observations clustered within the same article, recognizing that instances within an article may be shaped by shared context, author style, or topical focus. Each LLM-identified mention (instance) was coded using two binary variables:
  1. 1.
    Instance Type, where 1 = Individualized and 0 = Grouped.
  2. 2.
    Side, where 1 = Palestinian and 0 = Israeli.
To account for clustering, each instance was linked to a unique Article ID, ensuring that multiple mentions from the same article were recognized as non-independent.
We estimated four separate GEE logistic regression models, one for each media outlet (AJE, BBC, CNN, and NYT) In each model, the dependent variable was the binary Instance Type, and the independent variable was Side. We specified an exchangeable working correlation structure and used robust (Huber-White) standard errors to account for within-article correlation.
The estimated coefficients represent the log odds of an instance being individualized rather than grouped, comparing Palestinian to Israeli mentions. To aid interpretation, we also report odds ratios (exponentiated coefficients). A positive and statistically significant coefficient (or odds ratio >1>1) indicates that mentions of the Palestinian side are more likely to be individualized relative to Israeli mentions, while a negative coefficient (or odds ratio <1<1) indicates the reverse.
Measuring Narrative Elements Associated with Empathy in News Texts
To systematically assess the level of empathy conveyed in story narratives, we adopted a validated framework from Shen et al. (2024) [48]. In their study, the authors identified a set of narrative style elements that contribute to eliciting empathy and integrated them into the Human Empathy and Narrative Taxonomy (HEART). Each element in the taxonomy represents a distinct dimension of empathy, operationalized through targeted prompts for large language models (LLMs). For example, under the Vividness of Emotions metric, the LLM is instructed to assign a score of 1, 2, or 3 to indicate whether the emotional expressions in a text are not vivid at all, somewhat vivid, or extremely vivid.
Shen et al. (2024) found that three metrics contributed significantly to empathy:
  1. 1.
    Vividness of Emotions: the degree to which emotions are expressed vividly in stories.
  2. 2.
    Plot Volume: the frequency and significance of events in a story.
  3. 3.
    Character Development: the depth and evolution of characters throughout a narrative.
Of these, we retained only the first two. Character Development was excluded, as it is generally ill-suited to news reporting, where stories are concise and rarely allow for sustained character arcs, unlike the longer-form social media and podcast narratives used in the original study.
Given that the HEART prompts had already been validated against human-coded responses and shown to achieve “reasonable human-level” agreement, we adopted them with minimal adaptation. Specifically, we modified the in-prompt examples to better reflect the kinds of content the LLM would encounter in our corpus while preserving the conceptual intent of the originals. For instance, an example rated as 3 under Vividness of Emotions in Shen et al.:
“The pain of losing someone is like being stabbed in the chest. I was devastated when I lost her”
was replaced with:
“And you lived there in a horrific sense of fear. Every second that you live with this feeling is a terrible feeling, that you don’t really know if you’re going to wake up in the morning, or in a minute, if a missile is going to fall on you, if they’re going to come in with a Kalashnikov and start spraying us with bullets. The conditions are very, very difficult there.”
Similarly, an example rated as 1 for lack of vividness:
“I didn’t feel great about the situation”
was replaced with:
“Farhan al-Qadi, 52, a member of Israel’s Bedouin minority who was working at a kibbutz when he was seized in the Hamas raid of Oct 7.”
We also added a brief introductory statement in the prompt to orient the LLM to the type of text it would process. Other than these targeted adjustments, the structure, instructions, and scoring criteria from the original HEART prompts were preserved.
After scoring each story individually using these prompts, we addressed the fact that a single news article could contain multiple stories, often from both Israeli and Palestinian perspectives. To account for this, we aggregated the story-level HEART scores to create an article-level empathy metric, enabling us to capture the relative balance of empathy elicited for each side within the same piece of reporting.
For each of the two retained HEART metrics (Vividness of Emotions and Plot Volume), the calculation proceeded as follows:
  1. 1.
    Aggregate story-level ratings by side: For each article, we summed the ratings for all Palestinian stories to obtain a Palestinian aggregate score, and likewise summed the ratings for all Israeli stories to obtain an Israeli aggregate score.
  2. 2.
    Compute the difference: We subtracted the Israeli aggregate score from the Palestinian aggregate score. This yielded a single article-level empathy difference score for that metric, where positive values indicated higher empathy toward Palestinian stories, and negative values indicated higher empathy toward Israeli stories.
For example, consider an article containing three Palestinian stories with Vividness of Emotions ratings of 33,22, and $2$ (aggregate =7=7), and two Israeli stories with ratings of $1$ and $2$ (aggregate =3=3). The article-level empathy difference score for Vividness of Emotions would be: 73=47-3=4. A score of +4+4 in this case suggests that, on balance, the Palestinian narratives in the article were described with greater emotional vividness than the Israeli narratives. The same procedure was applied independently for Plot Volume, producing two article-level empathy difference scores, one per metric.
Baseline Model of Expected Weekly Casualty Reporting
To estimate the expected weekly reporting of casualty-related CVNs for each side across four news sources during the first year, we propose a baseline model. This model preserves the total weekly mentions for each media outlet, adjusts for significant events, and incorporates a decay factor to account for the natural decline in coverage over time.
The model consists of the following components:
  1. 1.
    Reporting Ratio: The reporting ratio $r_{n}(t)$ captures the overall tendency of news source $n$ to report casualty-related CVNs relative to actual casualty numbers in week $t$. It is defined as:
    $r_{n}(t)=\frac{M_{p,n}(t)+M_{i,n}(t)}{C_{p}(t)+C_{i}(t)}$
    where:
    • $M_{s,n}(t)$: the actual number of casualty-related CVN mentions in week $t$ by news source $n$ for side $s$, where $s\in\{p,i\}$, representing Palestine ($p$) and Israel ($i$)
    • $C_{s}(t)$: the actual number of casualties reported in week $t$ for side $s$
  2. 2.
    Spike Factor: The spike factor $S_{s}(t)$ accounts for abrupt increases in casualty numbers. It is defined as:
    $S_{s}(t)=\begin{cases}2,&\text{if }C_{s}(t)\geq 2\times C_{s}(t-1)\\ 1,&\text{otherwise}\end{cases}$
  3. 3.
    Initial Mentions: The initial mentions $G_{s,n}(t)$ generated in week $t$ for side $s$ by news source $n$ are given by:
    $G_{s,n}(t)=r_{n}(t)\times M_{s,n}(t)\times S_{s}(t)$
  4. 4.
    Cumulative Expected Mentions with Decay: The expected number of mentions $M^{\prime}_{s,m}(t)$ for side $s$ in week $t$ by news source $n$ incorporate contributions from initial mentions in the current and prior weeks, adjusted by side-specific decay factors $\delta_{p}=0.6$ (Palestine) and $\delta_{i}=0.8$ (Israel); see Side Note 1 below for the rationale behind the choice of side-specific decay factors:
    $M^{\prime}_{s,n}(t)=\sum_{k=0}^{t-1}G_{s,n}(k)\cdot\delta_{s}^{t-k}+G_{s,n}(t)$
    where:
    • $\delta_{s}$: the decay factor, where $\delta_{p}=0.6$ for $s=p$ (Palestine) and $\delta_{i}=0.8$ for $s=i$ (Israel)
    • $G_{s,n}(k)\cdot\delta_{s}^{t-k}$: the decayed mentions from previous week $k$.
    • $G_{s,n}(t)$:the initial mentions for the current week, which do not decay until subsequent weeks.
  5. 5.
    Normalization of Weekly Mentions: Finally, to normalize the expected mentions and reflect the actual number of mentions for each week, we define the normalized values for each side as follows:
    The normalized reporting ratio $w_{s}(t)$ for each side $s$ in week $t$ is calculated by:
    $w_{s}(t)=\frac{M^{\prime}_{s,n}(t)}{M^{\prime}_{p,n}(t)+M^{\prime}_{i,n}(t)}$
    Then, the final normalized expected mentions for side $s$ is recalculated as:
    $M^{\prime}_{s,n}(t)=(M_{p,n}(t)+M_{i,n}(t))\times w_{s}(t)$
    This normalization ensures that the final expected mentions for both Palestine and Israel are reflective of the actual number of mentions for each side, adjusted for the different decay rates. The model effectively accounts for variations in reporting tendencies, the impact of major events, and the natural decay of coverage over time.
Side Note 1: Rationale for Side-Specific Decay Factors
The choice of decay factors ($\delta_{p}=0.6$ for Palestine and $\delta_{i}=0.8$ for Israel) reflects differences in the frequency and timing of occurring casualties, as well as how media coverage responds to these dynamics. The Palestinian side experienced a continuous stream of casualties throughout the first year, causing media coverage to decay more rapidly as fresh incidents emerged weekly. This aligns with Pfefferbaum et al. [63], who observed that ongoing crises lead to shifting media focus and faster declines in coverage for earlier reports. Additionally, Friedman & Sutton [81] suggest that the continuous nature of such events accelerates the turnover of media attention.
In contrast, Israeli casualties were concentrated in the initial week, leading to a slower decay factor. The media’s prolonged interest in these early events stems from their novelty and dramatic impact, consistent with research by Downing et al. and Rohner & Frey [64, 82], which shows that significant initial events can sustain attention longer in the absence of ongoing crises. Thus, the differing decay factors capture these dynamics and are supported by literature on media responses to crises and the temporal aspects of news reporting.

### v2

@@
 We investigated how media outlets handled (de-)humanizing reporting, quantified the human costs of war, and reported on child victims.
+In order to establish bias, we compared media reporting, wherever possible, with objective standards such as actual casualty rates, as explained in greater detail below.
 This allowed us to uncover nuanced media biases across different sources.
+Importantly, with the term ”bias”, we are not assuming any form of intentionality or malevolence, but describing observable asymmetries in coverage that reflect underlying reporting tendencies.
+The possible reasons for the occurrence of such biases will be discussed in later sections.
 (De-)humanizing the Victims: Individualized vs Category-based Reporting
@@
 Instead, the reporting style of Western media and, to a lesser degree of AJE, resembles what is known as “false balance”, namely the tendency to create the impression that two sides are equivalent when, in reality, they are not (as will be explained in more detail in the discussion section).
+It is important to note, however, that one would not necessarily expect a perfect match between actual casualty rates and casualty reporting, indeed, such correspondence was absent in all outlets, including AJE.
+Some degree of over-reporting of the events of October 7th compared to the following twelve months would be expected on the basis of well-documented psychological mechanisms.
+First, the October 7th massacre was extremely salient, involving a large number of victims in a short time span.
+Research on the availability heuristic shows that a single-mass casualty event tends to attract disproportionate attention and to be remembered more vividly than a higher, but gradually accumulating, death tolls [30].
+Second, events that occur first in time or space tend to receive greater attention than those that follow (e.g., the ballot-order effect [31].
+Thus, to the extent that the October 7 attack is perceived as the “starting point” of the conflict, it is understandable that it received over-proportional coverage.
+Nevertheless, such cognitive and temporal biases cannot account for the extent to which reporting equated casualties on both sides.
+For instance, while the 38 Israeli children killed on October 7th might be expected to draw more attention than the 38 Palestinian children killed, on average, each day over the subsequent year, it would be implausible for the cumulative 14,000 Palestinian child deaths to receive equivalent attention as those 38 Israeli deaths.
 Sustaining Parity in the Face of Discrepancy
@@
 Israel) across four news sources.
-The results, summarized in Figure 3c, reveal that BBC overwhelmingly leads in using doubt-casting phrases, with 915 CVN mentions, 98% of which target Palestinians—averaging over two CVNs doubted daily during the first year.
+The results, summarized in Figure 3c, reveal that BBC overwhelmingly leads in using doubt-casting phrases, with 915 CVN mentions, 98% of which target Palestinians, averaging over two CVNs doubted daily during the first year.
 This far exceeds CNN (145), NYT (131), and AJE (72).
@@
 These findings underscore in particular BBC’s significant reliance on doubt-casting phrases, particularly when reporting on Palestinian casualties, compared to other sources.
+Doubting Palestinian sources is certainly justified, given the inherent difficulties in providing precise victim estimates during wartime.
+Statistics released by Palestinian authorities in Gaza were indeed frequently corrected and updated as new information became available.
+However, the same holds true for victim estimates on the Israeli side, where official statistics were revised repeatedly for up to six months after the events.
+Notably, skepticism toward Palestinian figures was often amplified by the practice of explicitly associating them with Hamas, an organization formally designated as terrorist group in the United States [77] and the United Kingdom [78], thereby highlighting their perceived lack of credibility.
+This rhetorical framing tends to obscure the fact that similar or identical casualty estimates were also reported and validated by independent researchers, international institutions, and humanitarian organizations (e.g., [79]).
 Children in the Narrative: Imbalances in CVN Reporting
@@
 Together, these reporting patterns insinuate a false premise of symmetry, where, in reality, the entity and prevalence of human suffering is blatantly asymmetrical.
-How did BBC, CCN, and NYT journalists manage to create the illusion of balance in the face of extreme disparities in military power and human suffering?
-The main strategy was to persistently refer back to the 7th of October massacre, even at a distance of months and in the absence of new events involving Israeli victims.
+How did BBC, CNN, and The NYT reporting convey an impression of (false) balance despite the extreme disparities in military power and human suffering?
+One contributing factor was the frequent return to the events of October 7th, even months later and in the absence of new incidents involving Israeli victims.
 Our analysis suggests that reiterations of October 7 were particularly pronounced during periods when significant events occurred on both sides, often overshadowing coverage of Palestinian suffering and diverting attention from the toll in Gaza.
-This strategy, which contradicts the common “agenda decay” phenomenon, not only creates an illusion of balance, but also reiterates the idea that the 7th of October was the starting point and “cause” of the subsequent events.
+This practice, which contradicts the common “agenda decay” phenomenon, not only creates an illusion of balance, but also reiterates the idea that the 7th of October was the starting point and “cause” of the subsequent events.
+Importantly, we do not intend to suggest that the reiteration of October 7th was part of a deliberate strategy pursued by the analyzed media, however, even such unintentional biases may have contributed to the neglect of Palestians’ victimhood in public perception.
+Biases can arise from common and often automatic cognitive processes.
+As previously discussed, a single mass-casualty event tends to attract disproportionate attention and to be remembered more vividly than a higher but gradually accumulating death toll.
+In addition, events that occur first in time or space typically receive greater attention than those that follow.
+Together, these mechanisms may partly explain why the October 7th massacre remained a persistent reference point in subsequent reporting.
+At the same time, perfect parity in coverage should not be expected, given the profound asymmetry in casualties between the two sides and the unequal number of conflict-related events that could be reported.
 This reporting pattern resembles what is generally referred to as “false balance” or “false equivalence”.

### v3

@@
 Abstract
-October 7th 2023 marked the start of a war against Gaza, which is considered one of the most devastating wars in modern history and has led to a stark attitudinal divide within and between countries.
-To investigate the role of media bias in reporting on this asymmetrical warfare, we analyzed over 14,000 news articles published during the first year of war in three Western (The New York Times, BBC, CNN) and one non-Western English-language outlets (Al Jazeera English).
-Exploring the media narratives concerning Israeli and Palestinian victims experiencing hardship, we found three systematic biases in Western media.
-1) Compared to Palestinian victims, represented mainly as undifferentiated collectives, Israeli victims were more likely to be portrayed as identifiable individual human beings.
-2) Despite the striking difference in all forms of hardship (casualties, displacement, etc.), Western journalists created a false balance, equating Israeli and Palestinian suffering, by persistently referring back to the 7th of October massacre, even in the absence of new events involving Israeli victims.
-3) When reporting on numbers of Palestinian (vs.
-Israeli) victims, journalists used language that casts doubt about the credibility of the information and the reputation of the source providing it, thereby selectively undermining the reader’s trust in the information regarding Palestinian suffering.
-Together, our analysis reveals a series of systematic journalistic biases in high-profile Western media that are absent or greatly reduced in Al Jazeera.
+October 7, 2023 marked the start of a war against Gaza, one of the most devastating conflicts in modern history, which quickly produced a stark global attitudinal divide.
+To examine the role of media bias in shaping public understanding of this asymmetrical war, we analyzed more than 14,000 news articles published during its first year across three major Western outlets (The New York Times, BBC, CNN) and one non-Western English-language outlet (Al Jazeera English).
+Focusing on media narratives surrounding Israeli and Palestinian victims, we identify three systematic biases in Western coverage: (1) Identifiable Victim Reporting: Israeli victims were substantially more likely to be depicted as identifiable individuals, whereas Palestinian victims were predominantly represented as undifferentiated collectives.
+(2) Equalization Bias: Despite the profound asymmetry in casualties, displacement, and other forms of suffering, Western reporting repeatedly invoked the October 7 attacks to equalize Israeli and Palestinian victimhood, even in the absence of new Israeli-casualty events.
+(3) One-sided Doubt Casting: Journalists disproportionately used language that casts doubt on the credibility of casualty figures and the reliability of sources when reporting Palestinian (vs.
+Israeli) victim counts, selectively undermining trust in information about Palestinian suffering.
+Across all three phenomena, these patterns were either absent or greatly attenuated in Al Jazeera English.
+Taken together, our analysis uncovers a coherent set of systematic biases in high-profile Western media coverage of the Gaza war, with implications for how global audiences come to understand and morally evaluate the conflict.
 Introduction
@@
 This was followed by an extensive war of Israel against Gaza that - already by the end of the first year - had produced an unprecedented humanitarian crisis [2], the destruction of over half of Gaza’s homes [3], the devastation of the school, university, and health systems, extreme deprivation, and the forcible displacement of almost the entire population [1].
-After one year of war, the death toll among Palestinians in Gaza amounted to over 42,000 people [4], the majority of whom women and children, although the true death toll may exceed 186,000 once indirect deaths are taken into account [5].
+After one year of war, the death toll among Palestinians in Gaza amounted to over 42,000 people [4], the majority of whom women and children.
+However, independent estimates suggest that the true death toll may greatly exceed these official statistics, both for violent mortality and when non-violent excess death rates are taken into account [5, 6]
 Over 97,000 people were injured [4], one fourth of whom with life-changing injuries such as amputations [7].
-Different from other conflicts (such as Ukraine), the population is trapped in Gaza due to the 17-year-old air, sea and land blockade.
+Different from other conflicts (such as Ukraine), the population was trapped in Gaza due to the 17-year-old air, sea and land blockade.
 The war has been particularly dire on children.
@@
 The current analysis, informed by psychological research and theory, focuses on subtle media biases in war journalism.
-Specifically, we examine three critical facets: (1) (De-)humanizing reporting, assessing to what extent victims are portrayed as individual human beings or as undifferentiated collectives.
-Thus, we look at the balance between group-level reports and individualized stories, which are more likely to evoke empathy.
-(2) Quantifying the human costs of war, exploring whether media reports present numbers in a meaningful and proportional manner, accurately reflecting the distribution of victims (and casualties more specifically) over the 12-month period examined.
-(3) Reporting on child victims, focusing on how children are individualized and featured in media coverage, as well as how their suffering is quantified.
+Specifically, we examine three potential journalistic biases that may have played a role in creating or sustaining attitudinal divides:
+  1. 1.
+    Identifiable victim reporting: War journalism has been criticized for its primary focus on frontline reporting, often neglecting the human consequences for the civilian population, and for children in particular [21].
+The media may either quantitatively give little space to civilian victims or report on them in a way that conceals their full human qualities (dehumanization), including their intellect, their culture, their morality, their feelings, and their individuality.
+One way in which journalists may downplay the civilian experience of war is by reporting group-level experiences rather than individual stories.
+Social scientists such as Slovic [22] and Schelling [23] have long observed that people often remain indifferent to mass atrocities such as war, mass murder, and genocide, yet are profoundly moved by individual stories.
+This psychological tendency is captured by what scholars call identifiable victim effect [24]: people feel more compassion toward single victims and are more likely to take action (for instance, through donations) when text or images describe the suffering of a single individual rather than that of larger groups.
+Compassion quickly fades as the number of victims increases [22], with the exception of very cohesive and interconnected small groups such as families [25].
+For instance, in a classical study by Kogut and Ritov [26], a single child in need of a costly life-saving treatment elicited greater distress and received greater donations than a group of sick children with identical needs.
+This phenomenon, mostly investigated in experimental work, is driven by different mechanisms, most notably that identifiable victims attract greater attention and evoke greater emotional reactions (such as sympathy or distress) than statistical victims or large groups [22].
+Also, by providing personalized information (such as age or name), as is typically done for single identifiable victims, narratives tend to be more vivid, which in turn induces a sense of familiarity and closeness.
+To examine the relative humanization of victims through individual stories, we therefore analyzed the degree to which Israeli and Palestinian civilian victims were portrayed through individualized (rather than collective) narratives.
+In the absence of bias, one would expect that the ratio of individualized to group-based narratives would be approximately the same for both sides.
+To the best of our knowledge, this is the first large-scale media analysis of individualizing vs.
+group-level reporting in the context of war (for a small-scale qualitative analysis on identifiable victim reporting, see [27]).
+  2. 2.
+    Equalizing bias: The second journalistic bias investigated here refers to the tendency to deny hierarchies in human suffering.
+In asymmetrical conflict situations, in which victims on one side greatly outnumber those on the other, one might expect media coverage to roughly mirror this difference, whereas stark deviations from this baseline may suggest that not all lives are equally news-worthy.
+We coined the term Equalization Bias to describe the tendency to dedicate equal media attention or space to two groups that, in reality, show very different levels of human suffering.
+For instance, giving the same coverage to Bosnian and Serbian victims in the Yugoslav war would constitute a case of equalizing bias.
+The equalization bias resembles, and was inspired by, the concept of false balance [28], an established journalistic bias, where two positions are presented as epistemically equivalent and receive equal space or attention (e.g.
+airtime) despite differences in actual validity and trustworthiness, such as climate skeptics and climate scientists or pro- and anti-vaccination proponents [28, 29, 30].
+Both, the false balance and the equalization bias create a false equivalence between non-comparable objects, but whereas false balance revolves around the question of validity of information, the equalization bias proposed here alludes to the allocation of narrative resources to different groups (e.g., number of stories are told, space devoted, etc) that is not proportionate of the actual scale of human suffering.
+Furthermore, given the novelty of the proposed bias, we complement our large-scale media analysis with a controlled experiment designed to identify the psychological mechanism that may give rise to this bias.
+  3. 3.
+    One-sided doubt-casting: The third bias refers to linguistic devices (including hedges such as “allegedly”) that signal that a reported fact is unconfirmed and, therefore, that the claim is uncertain (e.g., Bailey et al., 2014 [31]).
+Such statements imply a lack of verifiability or consensus and, hence, are likely to induce doubt in the reader’s mind.
+In the context of war or natural disasters, where actual victim numbers can not easily be established or independently verified, journalists necessarily rely on external sources, typically governmental agencies such as the police or the health ministries of the countries involved.
+Whereas doubt-casting language is generally considered a sign of journalistic integrity and may reflect high ethical standards in situations of uncertainty, it becomes problematic when applied in a one-sided manner.
+To uncover potential bias, we therefore investigate the degree to which doubt-casting is used to either undermine the credibility of the source of information (e.g., according to the Hamas-controlled health ministry) or to question the reliability of the victim numbers reported in an article (such as allegedly; reportedly; appear).
+Having outlined the three forms of journalistic bias, we next clarify how they are explored in the present study.
+Importantly, throughout this work we use the term “bias” to describe systematic asymmetries or deviations in media coverage relative to objective baselines, such as actual casualty rates.
+The concept is applied to all three phenomena discussed above, without implying any form of intentional distortion or malevolent intent on the part of journalists.
+Possible explanations for the emergence of such patterns are discussed in later sections.
+Our analyses focus on the human consequences of war for the civilian population (including children), that constitute the overwhelming majority of victims on both, the October 7 attack (69% of all victims) and in the subsequent war on Gaza (estimated to be above 80% among Palestinian victims [32]).
+To investigate the three proposed biases, we employ two complementary analytical approaches: one centered on narrative representation and the other on quantitative reporting.
+  • Narrative representation examines how victims are portrayed in media stories, whether as individualized human beings or as undifferentiated collectives, capturing degrees of (de-)humanization in reporting.
+  • Quantitative reporting investigates how numbers are used to convey the human costs of war, assessing whether the distribution of attention and reported casualties reflects the actual scale, the asymmetry of suffering across groups and whether such numbers are questioned.
+Together, these two analytical dimensions allow us to systematically assess how subtle reporting tendencies may shape perceptions of conflict and contribute to the broader attitudinal divides observed across the international media landscape.
 Results
-In this study, we analyze over 14,000 news articles published between October 7, 2023, and October 7, 2024, from The New York Times (NYT), BBC, CNN, and Al Jazeera English (AJE).
-We focus solely on the textual content, deliberately excluding other media formats like videos, podcasts and transcripts to maintain consistency in the analysis.
-For a detailed breakdown of the data composition and collection methodology, refer to the Materials and Methods’ Data subsection.
-As we were interested in subtle biases, we purposefully focused on highly professional news outlets, while excluding tabloids or other popular news sources where blatant partisanship could be expected.
+In this study, we analyze over 14,000 news articles published between October 7, 2023 and October 7, 2024 by four major international media outlets: The New York Times (NYT), BBC, CNN, and Al Jazeera English (AJE).
+To ensure analytical consistency and cross-outlet comparability, we restrict our corpus to textual news content and exclude videos, podcasts, and transcripts.
+A detailed description of the outlet selection criteria, temporal scope, and data collection procedures is provided in the Materials and Methods (Data Selection and Sample Construction subsection).
 We start our analysis by examining the data landscape presented in Figure 1, which illustrates the total article counts by source (Figure 1a) alongside the temporal distribution of article counts across sources (Figure 1b).
@@
 AJE, by contrast, maintained a relatively high publication rate, which remained comparably steady even after major events.
-With this data, we conducted our key analyses.
-We investigated how media outlets handled (de-)humanizing reporting, quantified the human costs of war, and reported on child victims.
-In order to establish bias, we compared media reporting, wherever possible, with objective standards such as actual casualty rates, as explained in greater detail below.
-This allowed us to uncover nuanced media biases across different sources.
-Importantly, with the term ”bias”, we are not assuming any form of intentionality or malevolence, but describing observable asymmetries in coverage that reflect underlying reporting tendencies.
-The possible reasons for the occurrence of such biases will be discussed in later sections.
-(De-)humanizing the Victims: Individualized vs Category-based Reporting
-War journalism has been criticized for its primary focus on frontline reporting, often neglecting the human consequences for the civilian population, and for children in particular [20].
-The media may either quantitatively give little space to civilian victims or report on them in a way that conceals their full human qualities (dehumanization), including their intellect, their culture, their morality, their feelings, and their individuality.
-One way in which journalists may downplay the civilian experience of war is by reporting group-level experiences rather than individual stories.
-Social scientists such as Slovic [21] and Schelling [22] have long observed that people often remain indifferent to mass atrocities such as war, mass murder, and genocide, yet are profoundly moved by individual stories.
-This psychological tendency is captured by what scholars call identifiable victim effect [23]: people feel more compassion toward single victims and are more likely to take action (for instance, through donations) when text or images describe the suffering of a single individual rather than that of larger groups.
-Compassion quickly fades as the number of victims increases [21], with the exception of very cohesive and interconnected small groups such as families [24].
-For instance, in a classical study by Kogut and Ritov [25], a single child in need of a costly life-saving treatment elicited greater distress and received greater donations than a group of sick children with identical needs.
-This phenomenon is driven by different mechanisms, most notably that identifiable victims attract greater attention and evoke greater emotional reactions (such as sympathy or distress) than statistical victims or large groups [21].
-Also, by providing personalized information (such as age or name), as is typically done for single identifiable victims, narratives tend to be more vivid, which in turn induces a sense of familiarity and closeness.
-Unsurprisingly, the identifiable victim effect is more powerful when the victim is not responsible for their plight [23].
-This is particularly relevant when reporting on children who, by definition, cannot be held responsible for an armed conflict or for the causes that led up to it.
-As such, by examining both the degree of individualization in war reporting and the specific attention devoted to children, our analysis seeks to uncover whether media narratives amplify or diminish the humanity of those most powerless in the face of war.
-We began our analysis by identifying all instances in which individuals or groups were mentioned in news articles across our four media sources.
+Using this dataset, we examined the three biases outlined in the Introduction.
+We began with narrative representation, focusing on the first two biases, identifiable victim reporting and equalizing bias, by asking whose stories are individualized, whose remain collective, and how casualty-related reporting reflects underlying asymmetries in victimization.
+Narrative Representation as an Indicator of Identifiable Victim and Equalizing Bias
+Narrative representation concerns how civilian victims are made visible (or remain obscured) in news coverage.
+Individualized stories tend to humanize victims and elicit empathy, whereas collective or statistical references do so to a much lesser extent.
+Examining these narrative choices enables us to identify two forms of bias: whether some victims are more likely to be individualized (identifiable victim bias) and whether narrative space is distributed in ways that obscure large disparities in suffering (equalizing bias).
+Identifiable Victim Bias: Individualized vs Collective Reporting
+We began by examining the identifiable victim bias, that is, the tendency to portray some victims as individualized persons while others are presented as undifferentiated collectives.
 To do this, we employed a large language model (LLM) to extract every reference to civilians, whether individual or collective, and classify each instance as pertaining to either the Palestinian or Israeli side.
@@
 0.09), though the disparity was less stark (see Supplementary Table 3).
-In contrast, AJE showed a more balanced pattern, with nearly equal ratios for both sides (0.09 for Palestinians vs.
+In contrast, AJE showed a more balanced pattern, with similar ratios for both sides (0.09 for Palestinians vs.
 0.07 for Israelis).
@@
 We formalized this asymmetry using a Generalized Estimating Equations (GEE) regression analysis (see Materials and Methods, subsection “Testing Individualization Bias with GEE Models”), which confirmed that Western outlets were significantly more likely to individualize the Israeli side in their reporting.
-Specifically, we find that the odds of a reported story being individualized are 13.6% lower for Palestinian instances compared to Israeli ones in the case of BBC, 27.0% lower in the case of CNN, and 38.9% lower in the case of NYT.
+Specifically, we find that the odds of a reported story being individualized are 14.7% lower for Palestinian instances compared to Israeli ones in the case of BBC, 28.7% lower in the case of CNN, and 41.3% lower in the case of NYT.
 No statistically significant difference was observed for AJE.
-Imbalance in Casualty Reporting
+Taken together, these results provide clear evidence of an identifiable victim bias in Western coverage, whereby Israeli civilians are more often presented as fully individuated persons, whereas Palestinians are more frequently portrayed as part of an anonymous collective.
+Equalizing Bias in Casualty Reporting
 We then turned our attention to the content of the stories themselves, categorizing each instance based on the most severe type of hardship described.
@@
 Among these, we opted to concentrate on stories labeled as “casualties” (i.e., deaths and injuries), as this category was the most frequent across all sources, relevant to both sides, and supported by reliable statistics from international organizations such as the UN, allowing for systematic comparison.
-When examining the ratio of Palestinian to Israeli casualty-related stories, as shown on the left side of Figure 2b, a stark contrast was revealed between AJE and Western outlets.
+After identifying casualty-related stories, we next assessed the presence of equalizing bias by comparing how frequently outlets reported on Palestinian versus Israeli casualties.
+As shown on the left side of Figure 2b, a stark contrast was revealed between AJE and Western outlets.
 AJE published roughly eight Palestinian casualty stories for every Israeli story, while BBC, CNN, and NYT maintained a close to one-to-one ratio.
@@
 This stark disparity in actual casualties reflects the marked power differential between the two groups, and contrasts with the seemingly “balanced” reporting in Western media.
-In theory, if all events were covered equally and without selection bias, one would expect a proportional discrepancy in the number of stories reflecting the underlying reality.
-Instead, the reporting style of Western media and, to a lesser degree of AJE, resembles what is known as “false balance”, namely the tendency to create the impression that two sides are equivalent when, in reality, they are not (as will be explained in more detail in the discussion section).
-It is important to note, however, that one would not necessarily expect a perfect match between actual casualty rates and casualty reporting, indeed, such correspondence was absent in all outlets, including AJE.
-Some degree of over-reporting of the events of October 7th compared to the following twelve months would be expected on the basis of well-documented psychological mechanisms.
+In theory, if all events were covered equally and without selection bias, one would expect a roughly proportional discrepancy in the number of stories reflecting the underlying reality.
+Instead, the reporting style of Western media and, to a lesser degree of AJE, is coherent with our phenomenon of equalizing bias, namely the tendency to create the impression that two sides are equivalent when, in reality, they are not.
+In particular, Western outlets create a false equivalence by giving equal narrative weight to Israeli experiences, despite the vastly different realities of suffering on the ground.
+It is important to note, however, that one would not necessarily expect casualty reporting to mirror casualty rates perfectly.
+In fact, no outlet, including AJE, displayed a fully proportional relationship.
+Some degree of over-reporting of the events of October 7th relative to the subsequent months is psychologically plausible.
 First, the October 7th massacre was extremely salient, involving a large number of victims in a short time span.
-Research on the availability heuristic shows that a single-mass casualty event tends to attract disproportionate attention and to be remembered more vividly than a higher, but gradually accumulating, death tolls [30].
-Second, events that occur first in time or space tend to receive greater attention than those that follow (e.g., the ballot-order effect [31].
-Thus, to the extent that the October 7 attack is perceived as the “starting point” of the conflict, it is understandable that it received over-proportional coverage.
-Nevertheless, such cognitive and temporal biases cannot account for the extent to which reporting equated casualties on both sides.
-For instance, while the 38 Israeli children killed on October 7th might be expected to draw more attention than the 38 Palestinian children killed, on average, each day over the subsequent year, it would be implausible for the cumulative 14,000 Palestinian child deaths to receive equivalent attention as those 38 Israeli deaths.
-Sustaining Parity in the Face of Discrepancy
+Research on the availability heuristic shows that single-mass casualty events tend to attract disproportionate attention and are remembered more vividly than larger, but gradually accumulating, death tolls [37].
+Second, events that occur first in time or space often receive greater attention than those that follow; a pattern observed in phenomena such as the ballot-order effect [38].
+Thus, if October 7 is perceived as the “starting point” of the war, some degree of disproportionate coverage would be understandable.
+Nevertheless, such cognitive and temporal biases cannot account for the extent to which media reporting equated casualties on both sides.
+For instance, while the 38 Israeli children killed on October 7th might be expected to draw more attention than 38 Palestinian children killed, on average, each day over the subsequent year, it would be implausible for the cumulative 14,000 Palestinian child deaths over the course of the year to warrant narrative attention equivalent to those 38 Israeli deaths.
+This pattern is consistent with equalizing bias: a structural tendency to allocate roughly equal narrative space to two sides despite profound asymmetries in underlying human suffering.
+Such patterns may be shaped, in part, by asymmetries in identification with conflict parties, which we next attempted to test experimentally.
+Experimental Evidence on Identification as a Mechanism of Equalizing Bias
+To examine whether equalizing bias can arise from social identification alone, i.e.
+the degree to which individuals perceive a group as part of their social self and feel attached to it, we conducted a controlled experiment in which participants in the US and UK were presented with a fictitious asymmetric geopolitical conflict between two countries differing in military strength and civilian death tolls (10 vs.
+40).
+Participants were randomly assigned to adopt the role of a journalist aligned with the stronger country, the weaker country, or an independent international outlet (see Supplementary Note 2 for full details).
+Participants reviewed brief profiles of 50 civilian casualties (10 from the stronger country; 40 from the weaker country) and were asked to select 10 victims whose stories they would feature in their news articles.
+In the absence of bias (random selection), proportional coverage would yield, on average, two strong-country victims and eight weak-country victims.
+Consistent with the equalizing bias observed in our media analysis, participants assigned to the stronger-country condition substantially over-selected strong-country victims (mean of 3.7 rather than 2) relative to the proportional benchmark.
+Participants assigned to the weaker-country condition did not show such over-selection (mean of 1.8 rather than 2), while independent participants occupied an intermediate position.
+Critically, identification predicted selection behavior: the more participants identified with the strong country, the more strong-country victims they selected ($r=.43$, $p<.001$).
+Further supporting the critical role of identification, a mediation model (see Supplementary Note 2) demonstrated that the equalizing bias is driven by journalists’ identification with the country to which they were assigned.
+Together, these findings provide causal evidence that equalizing bias can arise from asymmetries in social identification, lending support to the interpretation that the patterns observed in BBC, CNN, and NYT coverage are shaped by stronger identification with Israel.
+Sustaining Equalizing Bias in the Face of Discrepancy
 This led us to a critical line of inquiry: How do Western media outlets sustain parity in reporting casualty-related stories between both sides when the actual number of casualties is so drastically imbalanced?
@@
 This concentration persisted even in stories published several months after the attack.
-In other words, a single moment in time, the initial event, was repeatedly covered to sustain a consistent volume of Israeli casualty narratives, thereby flattening the temporal and numerical asymmetry in actual deaths.
+In other words, a single moment in time, the October 7 event, was repeatedly revisited to maintain a steady volume of Israeli casualty narratives, thereby flattening the temporal and numerical asymmetry in actual deaths to sustain equalizing bias.
 To explore this pattern further, we examined how Western media reported on the Israeli side during moments of significant Palestinian civilian casualties.
@@
 This suggests that the imbalance was not merely the result of coverage decay but was actively reinforced whenever Israeli stories could be interwoven with Palestinian casualty events, thereby sustaining a structural asymmetry in how loss and victimhood were represented.
-A more detailed explanation of our date-selection procedure and findings, including additional quotes, is provided in Supplementary Note 2.
+A more detailed explanation of our date-selection procedure and findings, including additional quotes, is provided in Supplementary Note 3.
 The number of Israeli stories, related to October 7, published by each outlet on the dates examined can be found in Supplementary Table 4.
-Quantifying Empathy-Related Narrative Elements
-To assess whether the qualitative imbalance we observed persisted across the full year of coverage, we quantified differences in narrative elements that can influence reader empathy, using the validated Human Empathy and Narrative Taxonomy (HEART) framework [50].
+Empathy-Related Narrative Elements in Individualized Stories
+Given that Israeli victims are more likely to be portrayed as identifiable individuals than Palestinian victims, the question arises whether the individualized stories about the former are also told in a more empathic narrative style.
+To explore this, we quantified differences in narrative elements that can influence reader empathy, using the validated Human Empathy and Narrative Taxonomy (HEART) framework [57].
 Specifically, we applied two HEART metrics: (i) Vividness of Emotions and (ii) Plot Volume (the main predictors of perceived empathy in the HEART model) to all hardship-related individualized stories in our corpus.
@@
 At the same time, it is worth noting that, on average, stories about Palestinians were written with greater emotional vividness across outlets (albeit less so for NYT), which is not surprising given the scale of Palestinian suffering.
-What is striking, however, is that even outlets such as BBC and CNN, while often writing about Palestinians in a more emotional way, still devoted disproportionate attention to Israeli narratives during key hostage-related events.
-Taken together, these findings suggest that the asymmetry lies less in the tone itself than in the balance of coverage: Western outlets often create a false equivalence by giving equal or greater narrative weight to Israeli experiences, despite the vastly different realities of suffering on the ground.
-False Balance in the Coverage of Child Victims
+Taken together, the above findings suggest that the asymmetry lies less in the narrative tone than in the relative frequency with which individualized stories are told in the first place: In Western media, Israeli suffering is more likely to be told through individual stories, whereas collective statements abound for Palestinian victims.
+Yet, once a journalist decides to report individualized stories about Palestinians, these stories are, on average, no less empathic towards the victims than those told about Israelis.
+Children as a test case of Equalizing Bias
+Children offer a particularly revealing test case for equalizing bias.
 Given the high vulnerability of children and the fact that they cannot be held responsible for armed conflict nor for the election of leaders who promote armed conflicts, it is not surprising that children have a special status in international humanitarian law (among which the UN Convention on the Rights of the Child [58]).
@@
 Children (including older children) are prioritized in helping decisions because they evoke greater nurturing concern and compassion [63], because they are generally not held responsible for their fate, and because they have, on average, a greater number of years ahead, which makes helping them a rational choice.
-As such, it is essential to understand how much attention the different media sources pay to child victims of war.
+As such, one might expect coverage to track the underlying disparity in child casualties especially closely.
 We therefore extended our analyses of individualized reporting to specifically track mentions of children.
@@
 Although they make up only 3% of Israeli deaths, their stories account for as much as 44% of child-related individualized reporting in CNN, with similarly elevated shares in BBC (36%) and NYT (40%).
-This disproportionate emphasis mirrors the false balance documented in our earlier analyses, where the relative weight given to Israeli and Palestinian suffering diverged sharply from the underlying reality.
+This disproportionate emphasis mirrors the equalization bias documented in our earlier analyses, where the relative weight given to Israeli and Palestinian suffering diverged sharply from the underlying reality.
 The imbalance becomes even clearer when looking at the distribution of child-related stories within each outlet (Supplementary Table 5b).
@@
 Instead, Western outlets devoted a strikingly large share of their child coverage to Israeli children: BBC 44%, CNN 43%, and NYT 51%.
-This pattern reflects a form of false balance: by the end of the first year of war, 38 Israeli [57] and 14,000 Palestinian children [7] had been killed, a disparity likely even greater for other hardships such as injuries, displacement, and famine.
+This pattern reflects a form of equalization bias: by the end of the first year of war, 38 Israeli [64] and 14,000 Palestinian children [8] had been killed, a disparity likely even greater for other hardships such as injuries, displacement, and famine.
 If casualty numbers alone were used as a baseline, individualized stories on Israeli children would be expected to account for less than 1% of all child references.
@@
 The most extreme deviation is observed in the New York Times, where Israeli and Palestinian children were given nearly equal attention, despite the enormous difference in the actual number of victims.
-Quantifying the Human Cost of War: the Importance of Numbers
-The reporting of victim numbers in news articles is crucial for conveying the magnitude of casualties and the impact of violence on communities.
+Quantitative Reporting as an Indicator of Equalizing Bias and One-Sided Doubt-Casting
+Quantitative reporting provides a second lens through which media bias may manifest.
+Whereas narrative representation concerns whose suffering is made visible through individualized storytelling, quantitative reporting examines how numerical information is allocated, sourced, and qualified across the two sides.
+This dimension allows us to assess two forms of bias: (i) equalizing bias, reflected in the distribution of casualty-related numbers that do not mirror the underlying asymmetry in victimization, and (ii) one-sided doubt-casting, reflected in selective source attribution or linguistic uncertainty applied disproportionately to one side’s figures.
+Having analyzed biases in narrative representation, we next assess how they manifest in the numerical reporting of the four news outlets.
+Because numerical information plays a central role in shaping how audiences understand the scale and distribution of harm, disparities in how numbers are presented offer an important window into these biases.
+Reporting of victim numbers in news articles is crucial for conveying the magnitude of casualties and the impact of violence on communities.
 Quantitative data not only provides readers with a clear understanding of the scale of incidents but also shapes public perception and policy responses to crime and victimization.
@@
 Recognizing the importance of reporting numbers in conflict scenarios, we employed an LLM to analyze all articles and extract instances containing numerical references to civilians being harmed in any capacity, such as “100 civilians were killed in that strike.” For clarity and convenience, these instances will be referred to as Civilian Victim Numbers (CVNs).
-A detailed explanation of how this information was extracted and validated can be found in Supplementary Note 3.
+A detailed explanation of how this information was extracted and validated can be found in Supplementary Note 4.
+These extracted CVNs allow us to examine whether numerical attention is allocated proportionally across the two sides (equalizing bias) and whether numerical claims are treated differently in terms of source attribution or skepticism (doubt-casting bias).
 We begin by examining the total number of articles containing at least one CVN.
@@
 However, in Western media specifically, Israeli CVNs still amount to more than half the number of Palestinian CVNs, despite the vast disparity in actual casualties.
-Once again, this points to a form of false balance, where Israeli suffering is elevated to a level of prominence in coverage that does not reflect the underlying reality.
-False Balance in Coverage: A Baseline Comparison of Casualty-Related Numbers
+Once again, this points to a form of equalizing bias, where Israeli suffering is elevated to a level of prominence in coverage that does not reflect the underlying reality.
+Equalizing Bias in Coverage: Baseline Comparison of Casualty-Related Numbers
 To assess whether the observed imbalance is limited to the early stages of the war following the October 7 attack, we analyzed the aggregate article counts from each news source over time.
@@
 For a detailed description, kindly refer to the subsection “Baseline Model of Expected Weekly Casualty Reporting” in Materials and Methods.
+By comparing expected versus observed numerical attention, the model allows us to quantify the degree of equalizing bias for each outlet.
 The comparison between the expected number of casualty-related CVN mentions, based on the baseline model, and the actual number of casualty-related CVN mentions for each side is visualized over time for each news source in Supplementary Figure 7.
 A consistent trend emerges across all outlets: Israeli CVNs appear to be overreported, while Palestinian CVNs seem to be underreported.
-However, comparing the magnitude of these gaps, between expected and actual mentions, across different news sources is challenging.
-Just how much overreporting or underreporting, and thus potential bias, exists relative to the baseline?
+However, comparing the magnitude of these gaps between expected and actual mentions across different news sources is challenging.
+Just how much overreporting or underreporting—and thus potential bias—exists relative to the baseline?
 To address this, we calculate the weekly proportion of expected mentions allocated to each side based on the baseline model for each news outlet.
@@
 Building on the previous observations, we conducted an analysis to investigate the possibility of selective citation of sources for CVNs.
-We utilized an LLM to evaluate each sentence containing a CVN, prompting it to determine whether a source was cited (for details, see Supplementary Note 4).
+We utilized an LLM to evaluate each sentence containing a CVN, prompting it to determine whether a source was cited (for details, see Supplementary Note 5).
 Subsequently, we examined the frequency of cited sources for CVNs on each side and compared patterns across different news outlets (examples of the most popular ways to cite sources for each side can be found in Supplementary Tables 7 and 8).
@@
 These disparities raise questions about potential biases in validating casualty figures based on the context.
-Creating Skepticism
-Next, we analyzed how referenced CVNs were phrased to determine whether they were presented in a “casting doubt” style (details can be found in Supplementary note 5).
+Creating Skepticism through Doubt-Casting
+Next, we analyzed how referenced CVNs were phrased to determine whether they were presented in a “casting doubt” style (details can be found in Supplementary Note 6).
 Using an LLM, we examined full sentences containing CVNs for phrases implying skepticism or questioning the source or the reported numbers.
@@
 Israel) across four news sources.
-The results, summarized in Figure 3c, reveal that BBC overwhelmingly leads in using doubt-casting phrases, with 915 CVN mentions, 98% of which target Palestinians, averaging over two CVNs doubted daily during the first year.
+The results, summarized in Figure 3c, reveal that BBC overwhelmingly leads in using doubt-casting phrases, with 951 CVN mentions, 98% of which target Palestinians, averaging over two CVNs doubted daily during the first year.
 This far exceeds CNN (176), NYT (166), and AJE (93).
 Common Source Doubting phrases like “Hamas-run health ministry” were heavily used by BBC, CNN, and NYT, whereas analogous expressions for Israel (such as “Knesset-run health ministry”) were absent.
-CNN and NYT also favored phrases like “Hamas-controlled enclave.” For Uncertainty in Numbers, terms such as “reportedly” and “claim” were predominantly used to describe Palestinian casualties by Western outlets, whereas AJE applied them more evenly across both sides (see Supplementary Tables 11–14 for detailed lists of phrases by outlet).
-Western sources attributed 71–90% of their cast-doubting phrases to doubting the sources of Palestinian CVNs, compared to 21% for AJE.
-Additionally, AJE was the only outlet to use Uncertainty in Numbers phrases in a balanced manner: 53% referred to Palestinians and 26% to Israelis, a proportion consistent with the fact that AJE published roughly twice as many articles about Palestinian CVNs as about Israeli ones.
-In contrast, the other outlets applied such language almost exclusively to Palestinians, with 3% or less referring to Israelis.
+CNN and NYT also favored phrases like “Hamas-controlled Gaza.” For Uncertainty in Numbers, terms such as “reportedly” , “allegedly”, and “claim” were predominantly used to describe Palestinian casualties by Western outlets, whereas AJE applied them more evenly across both sides (see Supplementary Tables 11-14 for detailed lists of phrases by outlet).
+Western sources attributed 77-89% of their cast-doubting phrases to doubting the sources of Palestinian CVNs, compared to 28% for AJE.
+Additionally, AJE was the only outlet to use Uncertainty in Numbers phrases in a balanced manner: 49% referred to Palestinians and 23% to Israelis, a proportion consistent with the fact that AJE published roughly twice as many articles about Palestinian CVNs as about Israeli ones.
+In contrast, the other outlets applied such language almost exclusively to Palestinians, with 2% or less referring to Israelis.
 These findings underscore in particular BBC’s significant reliance on doubt-casting phrases, particularly when reporting on Palestinian casualties, compared to other sources.
@@
 Children in the Narrative: Imbalances in CVN Reporting
-To conclude this section of our analysis, we examined how frequently media outlets referenced children in their reporting of civilian victim numbers (CVNs) for both Palestinians and Israelis (Supplementary Figures 6 and 7, and Supplementary Table 15).
+To conclude this section of our analysis, we examined how frequently media outlets referenced children in their reporting of civilian victim numbers (CVNs) for both Palestinians and Israelis (Supplementary Figures 8 and 9, and Supplementary Table 15).
 To contextualize these proportions, we benchmark them in Supplementary Table 15 against the approximate share of children among civilian deaths on each side: around 5% for Israeli civilians on October 7 (and roughly 3% when including soldiers, based on public compilation) [32, 64], compared to 40–44% for Palestinian civilians over the first year of the war (October 7, 2023–October 7, 2024) [86, 87].
@@
 These findings suggest that even highly professional (and relatively “liberal”) Western media tend to downplay the individuality of Palestinian victims experiencing the traumatic (and often fatal) consequences of war, while acknowledging, to a greater extent, the personal suffering of Israeli victims.
+This finding confirms what Grimm et al.
+(2025) [27] observed in a small-scale qualitative discourse analysis of 40 articles published in four German newspapers during the first 7 weeks of war, namely that Israeli victims tended to receive more personalized coverage than Palestinian victims.
 Based on previous literature on the identifiable victim effect [22], the Western reporting pattern is likely to elicit greater empathy and compassion for Israeli than for Palestinian victims.
@@
 Taken together, these findings suggest that it is not as much the style of writing than the under-reporting of individual stories that contributes to the dehumanization of Palestinians in Western media.
-The second systematic bias in Western media (and to a lesser degree in AJE) was the tendency to equate Israeli e Palestinian suffering despite the stark difference in all forms of hardship (casualties, displacement, arbitrary detention and kidnapping, destruction of the health, school, and university systems, etc.).
+The second systematic bias in Western media (and to a lesser degree in AJE) was the tendency to equate Israeli to Palestinian suffering despite the stark difference in all forms of hardship (casualties, displacement, arbitrary detention and kidnapping, destruction of the health, school, and university systems, etc.), We termed this pattern “equalizing bias”.
 In the absence of bias, news reports should reflect the actual gaps in human suffering.
 To give only one example, if Israeli child deaths are less than 1% and Palestinian children more than 99% of the total number of children killed during the first year of war, news reports would roughly reflect this discrepancy.
-In contrast to this assumption, we find a systematic deviation from actual discrepancies that emerges consistently across observations: (a) In Western media, 47% of individual narratives on children referred to Israeli children who constitutes less than 1% of the victims; (b) Including all age groups, the ratio of Palestinian to Israeli casualty-related individual stories was approximately 1:1 in Western media, despite an actual casualty ratio of 20:1.
-(c) Especially in Western media, Israeli Civilian Victim Numbers are over-reported and Palestinian Numbers under-reported compared to what would be expected based on the actual number of victims and the distribution over time.
+In contrast to this assumption, we find a systematic deviation from actual discrepancies that emerges consistently across observations: (a) In Western media, 47% of individual narratives on children referred to Israeli children who constitutes less than 1% of the victims; (b) Including all age groups, the ratio of Palestinian to Israeli casualty-related individual stories was approximately 1:1 in Western media, despite an actual casualty ratio of 20:1; (c) Especially in Western media, Israeli Civilian Victim Numbers are over-reported and Palestinian Numbers under-reported compared to what would be expected based on the actual number of victims and the distribution over time.
 Together, these reporting patterns insinuate a false premise of symmetry, where, in reality, the entity and prevalence of human suffering is blatantly asymmetrical.
+This raises the question of whether the equalizing bias is unique to the Israeli-Palestinian conflict or whether it constitutes a more general bias likely to occur whenever journalists (and possibly laypeople) identify with the dominant side of an asymmetrical conflict.
+What argues in favor of the latter interpretation is our experimental study replicating the equalizing bias among lay people acting as journalists in an imaginary, greatly tilted conflict.
 How did BBC, CNN, and the NYT reporting convey an impression of (false) balance despite the extreme disparities in military power and human suffering?
@@
 At the same time, perfect parity in coverage should not be expected, given the profound asymmetry in casualties between the two sides and the unequal number of conflict-related events that could be reported.
-This reporting pattern resembles what is generally referred to as “false balance” or “false equivalence”.
-False balance occurs when equal space or attention (e.g.
-airtime) is given to two sides that, in reality, differ in validity, such as climate skeptics and climate scientists [83] or pro- and anti-vaccination proponents [84, 71].
-Extending this concept to war journalism, false balance can manifest when unequal parties in a conflict are portrayed as equally culpable or their actions as equally justified.
-Regarding victims, a false balanced perspective may portray different sides involved in a conflict as equally affected by the conflict.
-This can obscure asymmetries in power, suffering, and in violations of international law, leading audiences to perceive the conflict as an intractable clash of opposing but equivalent sides.
-As one of the main consequences of false balance is the production of uncertainty [85], it is plausible that, in the context of war journalism, falsely balanced portrayals reduce public recognition of victimization or aggression, thereby dampening calls for accountability and action.
-Whether this applies to the present case remains an open question for future research.
-A third finding concerns the presumed reliability of the reported information and the trustworthiness of the source providing it.
+Yet these cognitive mechanisms alone can hardly explain the almost perfect parity in coverage we observe, particularly in light of the profound asymmetry in casualties and the greatly unequal number of conflict-related events that could be reported.
+We interpret the reporting pattern observed in Western media as an instance of what we have termed the equalizing bias, namely the tendency to portray two groups as equally affected by a conflict despite stark differences in suffering.
+Although this is beyond the scope of the present media analysis, the equalizing of unequal parties may not only obscure asymmetries in suffering, but also in power and in culpability (such as violations of international law), leading audiences to (mis-)perceive the conflict as an intractable clash of opposing but equivalent sides.
+As one of the main consequences of false equivalence is the production of uncertainty [88], it is plausible that, in the context of war journalism, equalizing portrayals will reduce public recognition of victimization and aggression, thereby dampening calls for accountability and action.
+Whether the equalizing bias does, indeed, have these effects remains an open question for future research.
+Furthermore, why would Western journalists portray Israeli suffering at par with Palestinian suffering?
+We can envisage at least three, not mutually exclusive reasons: First, the equalization bias may reflect the journalistic norm of pursuing balance and impartiality (although the same desire for balance was not visible in the choice of individuating and doubt-casting reporting).
+This interpretation, however, was not supported by our experimental simulation study in which the endorsement of journalistic norms failed to predict the magnitude of the equalizing bias.
+Second, to the degree that Western reporters may identify more with the Israeli than the Palestinian side [28], they may engage in what is known as “competitive victimhood”, namely the tendency to claim that one’s own group has suffered more than a relevant outgroup [89].
+In particular, competitive victimhood provides a tool for regaining moral ground whenever a group is accused of inflicting harm to the outgroup (Sullivan et al., 2012 [90]).
+This interpretation is supported by our experimental findings, which show that identification with the stronger country significantly predicted the equalizing bias and fully mediated the effect of experimental condition on victim selection, suggesting that group identification alone is sufficient to generate competitive victimhood dynamics.
+The third and most disquieting possibility is that the equalizing bias reflects deliberate journalistic malpractice involving editorial censorship and/or self-censorship [91].
+Regardless of which underlying mechanism drives the phenomenon and whether it occurs in an intentional or unconscious manner, the equalizing bias is likely to muddy public understanding of asymmetrical power relations and lead people to suspend judgement and action [28].
+Finally, our third finding concerns the presumed reliability of the reported information and the trustworthiness of the source providing it.
 Unlike AJE, Western journalists were found to engage in strategies aimed at planting seeds of doubt in their readership when reporting on Palestinian victims.
@@
 By discrediting sources in a subtle way, journalists undermine the reader’s trust in the information provided and in the reputation of the source.
-Together, the present media analysis has identified multiple biases in Western media that, together, may have contributed to the global divide in public opinion and to the delay in public outrage in the Western world in front of the unfolding genocide.
-Somewhat surprisingly, we did not discover a complementary pro-Palestinian bias in AJE, which - with a few exceptions, like the underreporting of Israeli children CVNs - displayed a relatively balanced reporting style.
+Together, the present media analysis identifies multiple biases in Western media that may have contributed to the global divide in public opinion and to the delayed public outrage in the Western world in response to the unfolding genocide.
+Somewhat surprisingly, we did not find a complementary pro-Palestinian bias in AJE, which—with a few exceptions, such as the underreporting of Israeli children’s CVNs—exhibited a relatively balanced reporting pattern.
+Scope and Generalizability
+Beyond the present case, the study contributes a theory-driven analytic framework for examining how media bias manifests in the coverage of asymmetrical warfare.
+This framework can be extended to compare media outlets, conflicts (e.g., Israel–Gaza vs.
+Serbia–Bosnia), or temporal phases within the same conflict (e.g., Gaza wars between 2008 and 2023).
+To exemplify the latter, we conducted a follow-up replication for our main bias measures to analyze how the same media (AJE, BBC, CNN, NYT) reported on the previous 3 wars on Gaza; for a detailed analysis see Supplementary Note 7.
+Just like the most recent (2023) war, these three wars, which took place in 2008, 2012, and 2014, showed a strong asymmetry in power and victimhood with Palestinians representing over 90% of all casualties.
+Importantly, during the 2008 and 2012 wars, Israel was accused of war crimes, but not genocide.
+In 2014, some organizations voiced genocide accusations, but these did not lead to formal court rulings, unlike the recent 2023 war, where such accusations were brought before the International Court of Justice.
+If accusations of ingroup wrong-doing increase the motivation to restore the lost moral ground through competitive victimhood (Noor et al., 2012 [92]; Sullivan et al., 2012[2] , Young & Sullivan, 2016[3] ) and if the equalizing bias offers a way to achieve this, then one would expect a greater equalizing bias the more serious the accusation of wrong-doing.
+A comparison across the four Gaza wars is broadly consistent with this reasoning.
+Equalization in casualty reporting appears to have been present in earlier wars and to have intensified over successive conflicts, becoming significantly most pronounced in 2023 (see Supplementary Figure 13), with the notable exception of AJE.
+It is also in line with another time trend observed in our analysis, namely the tendency to individualize Israeli (vs.
+Palestinian) victims more clearly in the more recent wars (2014 and 2023) where accusations became more serious (see Supplementary Figure 12).
+While this interpretation is post-hoc, emphasizing the victimhood of the more powerful group (Israel) remains relevant.
+As Noor et al.
+[92] argue, competitive victimhood serves several functions, including inflating the perceived threat from the out-group, justifying in-group violence, and mobilizing moral and material support from third parties.
+For our last analysis concerning doubt-casting language when reporting casualty statistics, both source-doubting (such as “Hamas-run health ministry”) and doubt-casting expressions (such as “reportedly”) were rare before 2023.
+They saw a steep increase concerning Palestinian casualty figures in 2023, specifically for BBC.
+Together, this timeline suggests an increase in Western media bias over time, which is most pronounced for the individualizing and the equalizing biases.
+This additional analysis illustrates that our framework lends itself to multiple comparisons, including historical trends.
 Limits and future directions
-Despite its extensive coverage, the present study has a number of limitations that should be acknowledged and that could serve as a starting point for future research.
+Despite its extensive coverage, the present study has several limitations that should be acknowledged and could serve as a starting point for future research.
 First of all, we took a conservative approach and included only highly respected and relatively “liberal” media in our analysis.
 This has almost certainly led to an underestimate of actual media bias.
-Second, our media analysis lacks a comparison to other conflict situations in order to understand whether the biases identified here are unique to Palestine and Israel.
-The Ukraine war may offer such a comparison standard, given the temporal proximity.
-Thus, future studies should not only broaden the scope to include more conservative media (such as Fox News or Daily Mail), but also extend the analysis to other conflict situations (such as Ukraine).
+Thus, future studies should broaden the scope to include more conservative media (such as Fox News or Daily Mail).
 Another important future development regards the extension of the analysis beyond the first year of war.
-It is conceivable that the pro-Israeli / anti-Palestinian bias in Western media may have declined during the second year of the ongoing genocide and that this may have contributed to the greater awareness and the shift in public opinion in the West.
-Investigating whether the media biases identified here did, indeed, shape public opinion, and to what degree, remains another important issue to be addressed in future research.
+It is conceivable that the pro-Israeli / anti-Palestinian bias in Western media may have declined during the second year of the ongoing genocide and that this may have contributed to greater awareness and the shift in public opinion in the West.
+Investigating whether the media biases identified here did indeed shape public opinion and to what extent remains another important issue to be addressed in future research.
 Only longitudinal or experimental methods will be able to address this question and allow inferences about the actual impact of the biases observed here on public opinion and collective action.
 Materials and Methods
-Data Collection
+Data Selection and Sample Construction
+Media Outlet Selection.
+Our analysis focuses on four major international news organizations: The New York Times (NYT), BBC, CNN, and Al Jazeera English (AJE).
+Outlet selection was guided by four criteria:
+  • publication in English,
+  • digital accessibility of archival content,
+  • the presence of independent reporting infrastructure (including field correspondents), and
+  • substantial global audience reach.
+Audience reach and influence rankings consistently identify BBC, NYT, and CNN among the most widely consumed English-language news organizations worldwide.
+According to PressGazette (2025) [93], BBC, NYT, and CNN rank first, second, and fourth, respectively, in global visits among the top 50 English-language news websites.
+MSN, which ranks third, was excluded because it functions as a news aggregator without its own reporting staff or field reporters.
+Among non-Western outlets, several Indian news websites (including India Times, Hindustan Times, The Indian Express, India Today, and One India) appear high in the rankings, but their readership is overwhelmingly domestic or diaspora-based.
+The highest-ranked truly international non-Western outlet is AJE (46th place).
+This pattern is consistent with rankings produced by PRlab [94], which list CNN, BBC, and NYT as the first, second, and third most influential outlets for news coverage, while AJE—ranked sixth—is the leading non-Western organization.
+The prominence of these outlets is further supported by audience research such as the 2024 Digital News Report [95].
+Using the share of respondents who reported seeing content from each brand in the past month as a measure of global reach, CNN (25%), BBC (22%), and NYT (12%) emerge as the top three news brands worldwide, while AJE (10%; ranked fifth) is again the highest among non-Western outlets.
+The audience scale is substantial: the three Western outlets alone attract an estimated average of 1.906 billion visits per month, including approximately 421 million unique visitors.
+Temporal Scope.
+We restrict our analysis to the first year of the war (October 7, 2023 - October 6, 2024), as this period captured both the overwhelming majority of global public attention and the phase of maximal media influence.
+Independent indicators of information demand, including Google Trends and Wikipedia pageview statistics (Supplementary Figures 1–2), show a sharp spike in interest immediately following October 7, followed by sustained levels of attention that remained substantially elevated throughout the first year relative to the subsequent period.
+This early attention surge marks the interval in which news organizations allocated exceptional resources, public reliance on media was greatest, and core interpretive frames crystallized.
+It is also the period containing most of the conflict’s humanitarian and military variance, when reporting was most intense and consequential for public understanding.
+By contrast, coverage beyond the first year occurs under conditions of markedly lower public attention and reduced reporting intensity.
+As audience demand stabilizes at lower levels, coverage becomes more routinized and episodic, and editorial investment contracts.
+Because media effects and framing dynamics are strongest during periods of high salience and interpretive uncertainty, the first year provides the most analytically informative window for assessing systematic reporting asymmetries.
+Examining this interval therefore allows us to evaluate patterns of representation under conditions of maximal scrutiny, when media influence and agenda-setting capacity are at their peak.
+Data Retrieval and Filtering.
 The New York Times (NYT), CNN, and Al Jazeera English (AJE) news articles were accessed via the LexisNexis Academic Web Services API (WSApi) [^1] LexisNexis NYU guide, https://guides.nyu.edu/lexisnexis-rest-api [^2] LexisNexis WSApi https://solutions.nexis.com/wsapi/, accessed through October 2024.
@@
 To account for clustering, each instance was linked to a unique Article ID, ensuring that multiple mentions from the same article were recognized as non-independent.
-We estimated four separate GEE logistic regression models, one for each media outlet (AJE, BBC, CNN, and NYT) In each model, the dependent variable was the binary Instance Type, and the independent variable was Side.
+We estimated four separate GEE logistic regression models, one for each media outlet (AJE, BBC, CNN, and NYT).
+In each model, the dependent variable was the binary Instance Type, and the independent variable was Side.
 We specified an exchangeable working correlation structure and used robust (Huber-White) standard errors to account for within-article correlation.